### v1

# The Hydra Map and Numen Formalisms for Collatz-Type Problems

Abstract.
This paper details a generalization of the formalism presented in the author’s 2024 paper, “The Collatz Conjecture & Non-Archimedean Spectral Theory: Part I — Arithmetic Dynamical Systems and Non-Archimedean Value Distribution Theory”, to the case of Hydra maps on the ring of integers $\mathcal{O}_{K}$ of a global field $K$. In addition to recounting these definitions, background material is presented for the necessary standard material in algebraic number theory and integration and Fourier analysis with respect to the $p$-adic Haar measure. This paper is meant to serve as a technical manual for use of Hydra maps and numens in future research.
Key words and phrases:
Collatz Conjecture; $3x+1$ map; $5x+1$ map; Numen; Hydra map; $p$-adic numbers; arithmetic dynamics; ultrametric analysis; $p$-adic analysis
1. Document Summary
As explained in the abstract, this paper is meant to be a foundational reference for works involving the class of generalized Collatz-type maps I call Hydra maps. Its primary purpose is to fix certain notations, definitions, conventions, and formalisms that are used throughout my works of research on this topic. Of particular import is the aim of giving a formalism which deals with both the archimedean and non-archimedean completions of the underlying number field in a unified way. Future versions of this paper will hopefully extend this unification to a fully rigorous adÚlic formulation.
Note 1.
Readers interested only in definitions may safely skip all proofs. Readers familiar with algebraic number theory may skip Section 3 entirely. Sections 6 & 7 contain the only genuinely novel analytic constructions. The definitions for the class of maps being studied are given in Section 5.
Section 2 contains basic notation, about 90% of which is completely standard. Non-standard notations include the author’s use of $\overset{a}{\equiv}$ to indicate congruence mod $a$ and his use of $\left[\cdot\right]_{a}$ to denote the standard representative element set of representatives of projections mod $a$. These notations, though minor, are crucial, as my formalism becomes unreadable without them.
Section 3 contains a refresher on the definitions and notations for seminorms and absolute values on rings and fields, and their use in defining the standard concept of a place as used in algebraic number theory, and using uniformizers to extend $p$-adic values to field extensions.
Section 4 contains standard, elementary definitions and results from the theory of dynamical systems, with the concept of a divergent trajectory being defined with respect to a given absolute value on a field.
Section 5 presents the definition of Hydra maps, the class of generalized Collatz-type map I am studying. This terminology is entirely non-standard, and therefore should not be skipped. Note: due to the rapidly evolving nature of this subject, earlier definitions of Hydras (such as those given in my dissertation [12]) may be inconsistent with the ones given here. The definitions given in this document will be used as the standard going forward. For arithmetical purposes, Hydras are defined here as maps on the ring of integers of a given global field, however, by linear algebra, they can be defined on vector spaces over $\mathbb{Q}$ or over a field of rational functions over a finite field. Moreover, the definition can certainly be generalized to Dedekind domains, though I will not do so in this version of this paper.
Section 6 introduces the string formalism (adapted from the theory of automata; see [1]) and uses it to construct the numen, $X_{H}$, of a given Hydra map $H$. The characterizing functional equation of $X_{H}$ is established, and Lemma 1 generalizes the work in [14] to show when and how $X_{H}$ may be extended from a function on $\mathbb{N}_{0}$ to a function on the ring of $p$-adic integers. The section concludes with a statement and sketched proof of the Correspondence Principle, one of the main findings of my earliest work [12, 14]. This relates the integral values taken by $X_{H}$ over $\mathbb{Z}_{p}$ to the periodic and/or divergent points of $H$.
Section 7 begins with a brief review of the $p$-adic Haar probability measure and its use in $p$-adic integration. Several basic $p$-adic integration formulae are introduced, along with the elementary theory of Pontryagin Duality from abstract harmonic analysis and its application to $p$-adic Fourier analysis. Novel content in this section includes Definition 20 on page 20, which gives the $\ell$-adic characteristic function of $X_{H}$, for varying places $\ell$ of the underlying global field $K$. The section also gives the definition of the $\ell$-adic Wiener algebras. These are the spaces of all functions $K_{\ell}\rightarrow\mathbb{C}$ expressible as the inverse Fourier transform of an $L^{1}$ function $K_{\ell}\rightarrow\mathbb{C}$, and provide a natural functional-analytic backdrop for giving a unified treatment of Fourier analysis across the various completions of $K$. Using the Wiener algebras, we can realize the $\ell$-adic characteristic function of $H$ as the Fourier transform of a self-similar Borel probability measure $d\mu_{H,\ell}$. (Equivalently, this measure is the probability distribution associated to the $\ell$-adic valued random variable induced by $X_{H}$.) The section concludes with some useful formulae for working with $X_{H}$ as an $\ell$-adic valued random variable in the case where $\ell$ is a non-trivial non-archimedean place of $K$.
2. Preliminary Notation
2.1. Notations for Sets
For any prime $p$, we write $\mathbb{Z}_{p}$ and $\mathbb{Q}_{p}$ to denote the ring of $p$-adic integers and fields of $p$-adic rational numbers, respectively, equipped with the standard $p$-adic absolute value $\left|\cdot\right|_{p}$, with $\left|p\right|_{p}=1/p$. $\mathbb{Z}_{p}^{\times}$ denotes the group of multiplicatively invertible $p$-adic integers; this is the set of all $p$-adic integers which are not congruent to 0 mod $p$. We write $\mathbb{C}_{p}$ to denote the field of $p$-adic complex numbers, i.e., the metric completion of the algebraic closure of $\mathbb{Q}_{p}$. Note that $\mathbb{C}_{p}$ is not spherically complete.
We write $v_{p}\left(\cdot\right)$ to denote the $p$-adic valuation, with $\left|\cdot\right|_{p}=p^{-v_{p}\left(\cdot\right)}$ and $v_{p}\left(0\right)\overset{\textrm{def}}{=}\infty$, where $\overset{\textrm{def}}{=}$ means “by definition”.
For any integer $N\geq 1$, we write $\mathbb{Z}/N\mathbb{Z}$ to denote the set $\left\{0,\ldots,N-1\right\}$. We write $\left(\mathbb{Z}/N\mathbb{Z}\right)^{\times}$ to denote the subset of $\left\{0,\ldots,N-1\right\}$ consisting of all integers co-prime to $N$. We will treat these sets as the ring of integers mod $N$ and the multiplicative group of units mod $N$, respectively, as the need arises.
[…]
For any real number $x$, we write $\mathbb{N}_{x}$ to denote the set of all integers $\geq x$ (thus, $\mathbb{N}_{0}=\left\{0,1,2,\ldots\right\}$; $\mathbb{N}_{1}=\left\{1,2,\ldots\right\}$, etc.). We write $\mathbb{Z}_{p}^{\prime}$ to denote $\mathbb{Z}_{p}\backslash\mathbb{N}_{0}$; note that this is the set of all $p$-adic integers with infinitely many non-zero $p$-adic digits. $\mathbb{P}$, meanwhile, denotes the set of prime numbers: $2,3,5,7,\ldots$.
Also, as the reader may have noticed, for aesthetic reasons, we write $p$-adic variables in lower-case $\mathfrak{fraktur}$ font.
2.2. Notation for Congruences
For any $\mathfrak{z}\in\mathbb{Z}_{p}$ and any integer $n\geq 0$, we write $\left[\mathfrak{z}\right]_{p^{n}}$ to denote the unique integer in $\mathbb{Z}/p^{n}\mathbb{Z}$ which is congruent to $\mathfrak{z}$ modulo $p^{n}$. For terminological purposes, we follow Yvette Amice in referring to the series representation $\sum_{n=-n_{0}}^{\infty}c_{n}p^{n}$ of a $p$-adic number $\mathfrak{y}\in\mathbb{Q}_{p}$ (be it rational or integral) as the Hensel series of $\mathfrak{y}$.
We adopt the convention of viewing $p$-adic integers as power series (and, likewise, $p$-adic rational numbers as Laurent series) in the variable $p$, and, importantly we write $p$-adic digits from left to right, in order of increasing powers of $p$. Thus, in $2$-adic notation, we would write:
$1$ $=\centerdot_{2}1$
$2$ $=\centerdot_{2}01$
$3$ $=\centerdot_{2}11$
$4$ $=\centerdot_{2}001$
$\vdots$
and so on. Here, we use $\centerdot_{2}$ as the “$2$-adic decimal point”. Throughout this series, all $2$-adic numbers written in base $2$ will have a $\centerdot_{2}$ in them, and, more generally, all $p$-adic numbers written in base $p$ will have a $\centerdot_{p}$ in them. For example:
(2.1) $11\centerdot_{2}011=\frac{1}{2^{2}}+\frac{1}{2}+2^{1}+2^{2}=6+\frac{3}{4}$
(2.2) $121\centerdot_{3}01=\frac{1}{3^{3}}+\frac{2}{3^{2}}+\frac{1}{3}+3^{1}=\frac{16}{27}+3$
We employ a non-standard notation for congruences:
$x$ $\overset{a}{\equiv}y$
(2.3) $\Updownarrow$
$x$ $=y\mod a$
To give some examples, for $\mathfrak{z}\in\mathbb{Z}_{p}$, $\mathfrak{z}\overset{p^{n}}{\equiv}k$ means $\mathfrak{z}$ is congruent to $k$ mod $p^{n}$”; i.e., $\mathfrak{z}\in k+p^{n}\mathbb{Z}_{p}$. Given $\mathfrak{x},\mathfrak{y}\in\mathbb{Q}_{p}$, we write $\mathfrak{x}\overset{1}{\equiv}\mathfrak{y}$ to mean $\mathfrak{x}-\mathfrak{y}=0\mod\mathbb{Z}_{p}$. In particular, note that:
  • $\mathfrak{y}\overset{1}{\equiv}\left\{\mathfrak{y}\right\}_{p}$ for all $\mathfrak{y}\in\mathbb{Q}_{p}$;
  • $s,t\in\hat{\mathbb{Z}}_{p}$ denote the same element of $\hat{\mathbb{Z}}_{p}$ if and only if $s\overset{1}{\equiv}t$;
  • $\mathfrak{z}\overset{1}{\equiv}k$ is true for all $\mathfrak{z}\in\mathbb{Z}_{p}$ and all $k\in\mathbb{Z}$.
For a rational integer $n$, $n\overset{2}{\equiv}0$ means $n$ is even, while $n\overset{2}{\equiv}1$ means $n$ is odd. Naturally, we write $x\overset{a}{\cancel{\equiv}}y$ to mean “$x$ is not congruent to $y$ mod aa”.
2.3. Notation for Convergence
[…]
Also, given a topological ring $R$, we write $\overset{R}{=}$ to mean that convergence occurs in the topology of $K$, where $K$ is the field of fractions of $R$, and that the terms of the sequence being limited or the partial sums of the series being limited are also elements of $R$. Thus, for a function $f:\mathbb{Z}_{2}\rightarrow\mathbb{Z}_{q}$, we write:
(2.4) $\lim_{n\rightarrow\infty}f\left(\left[\mathfrak{z}\right]_{2^{n}}\right)\overset{\mathbb{Z}_{q}}{=}f\left(\mathfrak{z}\right),\textrm{ }\forall\mathfrak{z}\in\mathbb{Z}_{2}$
to mean that, for each $\mathfrak{z}\in\mathbb{Z}_{2}$, the sequence $\left\{f\left(\left[\mathfrak{z}\right]_{2^{n}}\right)\right\}_{n\geq 0}$ consists of elements of $\mathbb{Z}_{q}$ that converge in $\mathbb{Z}_{q}$ to $f\left(\mathfrak{z}\right)$ as $n\rightarrow\infty$.
2.4. Preliminaries from $p$-Adic Analysis
For a refresher on the topic, the reader can refer to Robert’s book [8], Schikhof’s book [10], or the author’s doctoral dissertation [12]. For our immediate purposes, we need only a few results from the theory of ultrametric analysis:
Proposition 1.
The $p$-adic absolute value $\left|\cdot\right|_{p}$ satisfies the ultrametric inequality:
(2.5) $\left|\mathfrak{a}+\mathfrak{b}\right|_{p}\leq\max\left\{\left|\mathfrak{a}\right|_{p},\left|\mathfrak{b}\right|_{p}\right\},\textrm{ }\forall\mathfrak{a},\mathfrak{b}\in\mathbb{Q}_{p}$
with equality if and only if $\left|\mathfrak{a}\right|_{p}\neq\left|\mathfrak{b}\right|_{p}$.
Proposition 2.
Let $\left(\mathbb{K},\left|\cdot\right|_{\mathbb{K}}\right)$ be a metrically complete non-archimedean field, and let $\left\{a_{n}\right\}_{n\geq 0}$ be a sequence in $\mathbb{K}$. Then, the series $\sum_{n=0}^{\infty}a_{n}$ converges in $\mathbb{K}$ if and only if $\lim_{n\rightarrow\infty}\left|a_{n}\right|_{\mathbb{K}}\overset{\mathbb{R}}{=}0$.
Proposition 3.
Let $p$ be a prime number. Then, a $p$-adic number $\mathfrak{y}\in\mathbb{Q}_{p}$ is a rational number (i.e., an element of $\mathbb{Q}$) if and only if the sequence of $\mathfrak{y}$’s $p$-adic digits is eventually periodic.
3. Global Fields, Absolute Values, & Completions
By a base field $\mathbb{F}$, we mean either the field of rational numbers ($\mathbb{Q}$), or the field $\mathbb{F}_{q_{\mathbb{F}}}\left(T\right)$ of formal rational functions of $T$ with coefficients in $\mathbb{F}_{q_{\mathbb{F}}}$, the finite field of characteristic (char) $p_{\mathbb{F}}$ and order $q_{\mathbb{F}}=p_{\mathbb{F}}^{n_{\mathbb{F}}}$, where $p_{\mathbb{F}}$ is a prime number and $n_{\mathbb{F}}$ is a positive integer. The ring of integers of $\mathbb{F}$, denoted $\mathcal{O}_{\mathbb{F}}$, is either $\mathbb{Z}$ (if $\mathbb{F}=\mathbb{Q}$) or $\mathbb{F}_{q_{\mathbb{F}}}\left[T\right]$, the ring of formal polynomials in $T$ with coefficients in $\mathbb{F}_{q_{\mathbb{F}}}$ (if $\mathbb{F}=\mathbb{F}_{q_{\mathbb{F}}}\left(T\right)$). By a global field $K$, we mean a finite-degree extension of $\mathbb{F}$. The ring of integers of $K$ is denoted $\mathcal{O}_{K}$. We write $d$ (or $d_{K}$, for emphasis) to denote the degree of $K$ over $\mathbb{F}$.
Given a prime ideal $\mathfrak{Q}$ of $\mathcal{O}_{K}$, recall that there exists a unique prime ideal $\mathfrak{P}$ of $\mathcal{O}_{\mathbb{F}}$ so that
(3.1) $\mathfrak{Q}\cap\mathcal{O}_{\mathbb{F}}=\mathfrak{P}$
and we then say that $\mathfrak{Q}$ lies over $\mathfrak{P}$, and that $\mathfrak{P}$ lies under $\mathfrak{Q}$.
Definition 1.
A seminorm on $R$ is a function $\left\|\cdot\right\|:R\rightarrow\left[0,\infty\right)$ so that:
(3.2) $\left\|0_{R}\right\|$ $=0$
(3.3) $\left\|1_{R}\right\|$ $=1$
(3.4) $\left\|r+s\right\|$ $\leq\left\|r\right\|+\left\|s\right\|,\textrm{ }\forall r,s\in R$
(3.5) $\left\|rs\right\|$ $\leq\left\|r\right\|\left\|s\right\|,\textrm{ }\forall r,s\in R$
If, in addition, $\left\|\cdot\right\|$ satisfies the ultrametric inequality:
(3.6) $\left\|f+g\right\|\leq\max\left\{\left\|f\right\|,\left\|g\right\|\right\}$
with equality whenever $\left\|f\right\|\neq\left\|g\right\|$, we say $\left\|\cdot\right\|$ is non-archimedean. We say $\left\|\cdot\right\|$ is multiplicative if $\left\|rs\right\|=\left\|r\right\|\left\|s\right\|$ for all $r,s\in R$.
We call $\left\|\cdot\right\|$ a ring norm if $\left\|r\right\|=0$ if and only if $r=0_{R}$. A ring absolute value is a ring norm which is multiplicative.
A valued ring is a ring with an absolute value. A normed ring is a ring with a norm. A completed valued ring is a valued ring so that the metric induced by the absolute value is complete. A Banach ring is a normed ring so that the metric induced by the norm is complete. A valued field is a valued ring which is also a field.
A normed ($R$-)algebra is an $R$-algebra $\mathcal{A}$ which, as a ring, is a normed ring, and which satisfies $\left\|r\alpha\right\|=\left\|r\right\|\left\|\alpha\right\|$ for all $r\in R$ and all $\alpha\in\mathcal{A}$. A Banach ($R$-)algebra is a normed algebra so that the metric induced by the norm is complete.
By the quality of an absolute value, field, norm, normed, space, Banach algebra, or the like, we refer to the property of whether the thing under consideration is archimedean or non-archimedean. Thus, we can have a seminorm of archimedean quality, a valued ring of non-archimedean quality, and so on.
Finally, two absolute values $\left|\cdot\right|_{a}$ and $\left|\cdot\right|_{b}$ are said to be equivalent if there is a real number $c>0$ so that:
(3.7) $\left|r\right|_{b}=\left|r\right|_{a}^{c},\textrm{ }\forall r\in R$
An absolute value is trivial if $\left|r\right|=1$ for all $r\in R\backslash\left\{0\right\}$.
Definition 2.
A place of a ring or field is an equivalence class of equivalent absolute values. Given a place $\ell$ of the field $K$, we then write $K_{\ell}$ to denote the metric completion of $K$ with respect to any absolute value in $\ell$; up to isomorphism, $K_{\ell}$ is independent of the choice of the representative absolute value in $\ell$. Places whose absolute values are non-archimedean are called finite places; places whose absolute values are archimedean are called infinite places.We write $\textrm{Pl}\left(K\right)$ to denote the set of all places of $K$, and write $\textrm{Pl}^{\times}\left(K\right)$ to denote the set of all non-trivial places of $K$. We write $\textrm{Pl}^{\infty}\left(K\right)$ to denote the set of all non-trivial archimedean places of $K$, and write $\textrm{Pl}^{0}\left(K\right)$ to denote the set of all non-trivial non-archimedean places of $K$. Recall that non-archimedean places are in a bijective correspondence with the prime ideals of $\mathcal{O}_{K}$, while archimedean places are in a bijective correspondence with the embeddings of $K$ in $\mathbb{C}$. If $\textrm{char}\left(K\right)>0$, $K$ has no archimedean places.
Places will be denoted by $\ell$. $K_{\ell}$ denote the completion of $K$ with respect to absolute values in $\ell$, respectively. The symbol $\infty$ will be used as a subscript to indicate an archimedean place. Thus, just as $\mathbb{Q}_{p}$ is the $p$-adic completion of $\mathbb{Q}$, $\mathbb{Q}_{\infty}$ is the archimedean completion of $\mathbb{Q}$, otherwise known as $\mathbb{R}$. In an abuse of notation, given a prime ideal $\mathfrak{P}\subset\mathcal{O}_{K}$, we will sometimes write $\left|\cdot\right|_{\mathfrak{P}}$ to denote an absolute value of the corresponding place. Given a non-archimedean place $\ell$ of $K$, we write $v_{\ell}$ to denote the associated valuation (note: valuation, not absolute value). If $\mathfrak{P}$ is the prime ideal associated to $\ell$, we also write $v_{\mathfrak{P}}$ to denote $v_{\ell}$.
Given places $\ell$ of $K$ and $q$ of $\mathbb{F}$, recall that we say $\ell$ lies over $q$ whenever $\ell$ and $q$ are of the same quality, and whenever:
• If $\ell$ and $q$ are non-archimedean, the prime ideal of $\mathcal{O}_{K}$ associated to $\ell$ lies over the prime ideal of $\mathcal{O}_{\mathbb{F}}$ associated to $q$;
• If $\ell$ and $q$ are archimedean, $q$ is the standard archimedean absolute value on $\mathbb{Q}$.
Recall that the completion of a global field with respect to a non-trivial place is called a local field. If the place $\ell$ is non-archimedean, $\mathcal{O}_{K_{\ell}}$, the ring of integers of $K_{\ell}$, is the completion of $\mathcal{O}_{K}$ with respect to $\ell$, and is a local ring. In particular, there exists an element $\pi_{\ell}\in\mathcal{O}_{K_{\ell}}$ called the ($\ell$-adic) uniformizer so that $\pi_{\ell}\mathcal{O}_{K_{\ell}}$ is the unique maximal ideal of $\mathcal{O}_{K_{\ell}}$. If $\mathfrak{P}$ is a prime ideal of $\mathcal{O}_{\mathbb{F}}$ and $p$ is a prime number so that a uniformizer $\pi_{\mathfrak{P}}$ of $\mathcal{O}_{\mathbb{F}_{\mathfrak{P}}}$ (the ring of integers of the $\mathfrak{P}$-adic completion of $\mathbb{F}$) satisfies $\left|\pi_{\mathfrak{P}}\right|_{\mathfrak{P}}=p^{-1}$, recall there is then there a unique integer $e_{K,\ell}\geq 1$, called the ramification index of $\ell$ over $\mathfrak{P}$, so that:
(3.8) $\left|\pi_{\ell}\right|_{\ell}=p^{-1/e_{K,\ell}}$
Moreover, we have that $\left|\cdot\right|_{\mathfrak{P}}$ extends uniquely to $K_{\ell}$, with:
(3.9) $\left|\pi_{\ell}\right|_{\mathfrak{P}}=p^{-1/e_{K,\ell}}$
Definition 3.
We write $\textrm{End}_{\mathbb{F}}\left(K\right)$ to denote the $\mathbb{F}$-algebra of all $\mathbb{F}$-linear maps $K\rightarrow K$; we can view these as $d\times d$ matrices with entries in $\mathbb{F}$. We write $\textrm{End}_{\mathbb{F}}^{\times}\left(K\right)$ to denote the group of invertible $\mathbb{F}$-linear maps $K\rightarrow K$; this is a group under map composition.
Definition 4.
Given a place $\ell$ of $K$ and an $r\in\textrm{End}_{\mathbb{F}}\left(K\right)$, we write $\left\|r\right\|_{\ell}$ to denote the $\ell$-adic operator norm of $r$, defined by:
(3.10) $\left\|r\right\|_{\ell}\overset{\textrm{def}}{=}\sup_{\begin{array}[]{c}z\in K\\ \left|z\right|_{\ell}\leq 1\end{array}}\frac{\left|rz\right|_{\ell}}{\left|z\right|_{\ell}}$
so that $\left|rz\right|_{\ell}\leq\left\|r\right\|_{\ell}\left|z\right|_{\ell}$ for all $z\in K$.
Remark 1.
Since $K$ is dense in the metric completion $K_{\ell}$, observe that any $r\in\textrm{End}_{\mathbb{F}}K$ uniquely extends to a map $K_{\ell}\rightarrow K_{\ell}$ possessing the same $\ell$-adic norm as $r$.
Definition 5.
Given a non-zero prime ideal $\mathfrak{P}\subseteq\mathcal{O}_{K}$, we say $u\in\textrm{End}_{\mathbb{F}}^{\times}\left(K\right)$ is a $\mathfrak{P}$-adic isometry whenever:
(3.11) $\left|uz\right|_{\mathfrak{P}}=\left|z\right|_{\mathfrak{P}},\textrm{ }\forall z\in K$
Remark 2.
Since $K$ is dense in the metric completion $K_{\mathfrak{P}}$, observe that any $u\in\textrm{End}_{\mathbb{F}}^{\times}\left(K\right)$ which is a $\mathfrak{P}$-adic isometry then uniquely extends to a map $u:K_{\mathfrak{P}}\rightarrow K_{\mathfrak{P}}$, which is an isometry of $K_{\mathfrak{P}}$.
Proposition 4 ($\mu u$ Factorization (Non-Archimedean)).
Let $r\in\textrm{End}_{\mathbb{F}}^{\times}\left(K\right)$, and let $P=\left\{\mathfrak{P}_{1},\ldots,\mathfrak{P}_{n}\right\}$ be a collection of finitely many non-zero prime ideals of $\mathcal{O}_{K}$, with associated valuations $v_{\mathfrak{P}_{1}},\ldots,v_{\mathfrak{P}_{n}}$ and normalized non-archimedean absolute values $\left|\cdot\right|_{\mathfrak{P}_{1}},\ldots,\left|\cdot\right|_{\mathfrak{P}_{n}}$. Then, there exist a scalar $\mu\in K^{\times}$ and a $u\in\textrm{End}_{\mathbb{F}}^{\times}\left(K\right)$ so that $r$ decomposes in $\textrm{End}_{\mathbb{F}}^{\times}\left(K\right)$ as:
(3.12) $r=\varpi u$
where we identify $\varpi$ with the map $z\mapsto\varpi z$ in $\textrm{End}_{\mathbb{F}}^{\times}\left(K\right)$, and where $u$ is a $\mathfrak{P}_{m}$-adic isometry for all $m\in\left\{1,\ldots,n\right\}$
(3.13) $\left|uz\right|_{\mathfrak{P}_{m}}=\left|z\right|_{\mathfrak{P}_{m}},\textrm{ }\forall z\in K$
Moreover, as constructed, we have that:
(3.14) $\left|rz\right|_{\mathfrak{P}_{m}}$ $=\left|\varpi z\right|_{\mathfrak{P}_{m}}$
for all $z\in K$ and all $m\in\left\{1,\ldots,n\right\}$.
We call (3.12) the decomposition of $r$ over $P$. Given an ideal $\Lambda$ of $\mathcal{O}_{K}$, we call (3.12) the decomposition of $r$ over $\Lambda$ when $P$ is the set of prime ideals in the factorization of $\Lambda$.
4. Dynamical Systems
Note 2.
[2] is an easy introduction to the subject.
Fix a set $X\rightarrow X$ and a map $T:X\rightarrow X$.
Notation 1.
For any integer $k\geq 0$, we write $T^{\circ k}$ to denote the composition of $k$ copies of $T$, with $T^{\circ 0}$ being defined as the identity map.
Definition 6.
Recall that a periodic point of $T$ is an $x\in X$ so that $T^{\circ k}\left(x\right)=x$ for some $k\geq 1$. We say $x$ is pre-periodic if there exists a $k\geq 1$ so that $T^{\circ k}\left(x\right)$ is a periodic point of $T$. We say $x$ is a strictly pre-periodic point if $x$ is a pre-periodic point of $T$ which is not a periodic point of $T$. The forward orbit (sometimes called the trajectory) of $x$ under $T$ is the sequence:
(4.1) $\left\{T^{\circ k}\left(x\right)\right\}_{k\geq 0}$
A cycle $\Omega$ of $T$ is the forward orbit of a periodic point of $T$.
If $X$ is a valued field with absolute value $\left|\cdot\right|_{X}$, we say $x\in X$ is a divergent point of $T$ if :
(4.2) $\lim_{k\rightarrow\infty}\left|T^{\circ k}\left(x\right)\right|_{X}\overset{\mathbb{R}}{=}+\infty$
We use the phrase “a divergent trajectory of TT” to refer to the forward orbit of a divergent point (which may or may not exist, depending on $T$).
Next, we have orbit classes and their associated definitions
Definition 7.
Recall that an orbit class of $T$ is a non-empty set $V\subseteq X$ so that the pre-image of $V$ under $T$ is equal to $V$: $T^{-1}\left(V\right)=V$.
An orbit class $V$ of $T$ is said to irreducible if it cannot be written as $V=V_{1}\cup V_{2}$, where $V_{1}$ and $V_{2}$ are orbit classes of $T$ with $V_{1}\cap V_{2}=\varnothing$.
A backward orbit of some $x\in X$ is a sequence $\left\{x_{n}\right\}_{n\geq 1}$ in $X$ so that $x_{n+1}\in T^{-1}\left(\left\{x_{n}\right\}\right)$ for all $n\geq 0$. Note that if $T$ is not injective, a given $x_{0}\in X$ can have more than one backward orbit.
We say an orbit class $V$ is periodic if it is irreducible and contains a cycle. We say an orbit class $V$ is divergent if it is irreducible and contains a divergent point.
Lastly, we recall two elementary results from the theory of discrete dynamical systems. For proofs, see [14].
Proposition 5.
Let $T:X\rightarrow X$ be any map. Then, the collection of irreducible orbit classes of $T$ form a partition of $X$ into pair-wise disjoint sets. These sets are the equivalence classes of $X$ under the relation $\sim$ defined by:
(4.3) $x\sim y\Leftrightarrow\exists m,n\geq 0:T^{\circ m}\left(x\right)=T^{\circ n}\left(y\right)$
In particular, if $X$ is countable, then $X$ consists of at most countably many irreducible orbit classes of $T$.
Proposition 6.
Let $X$ be a valued field with absolute value $\left|\cdot\right|_{X}$, and let $Y$ be a discrete subset of $X$, in the sense that there is a $\delta>0$ so that:
(4.4) $\inf_{\begin{array}[]{c}x,y\in Y\\ x\neq y\end{array}}\left|x-y\right|_{X}\geq\delta$
Then, for any map $T:Y\rightarrow Y$ be any map, every irreducible orbit class of $T$ in $Y$ is either periodic or divergent. Consequently, any $x\in Y$ is either a divergent point of $T$ or a pre-periodic point of $T$.
Proof: Let $V$ be an irreducible orbit class of $T$ in $Y$.
If $V$ contains a cycle $\Omega$ of $T$, then, given any $x\in V$, $x\sim\omega$ for some periodic point $\omega\in\Omega$. So, $x$ is eventually iterated to some element of $\omega$’s trajectory. Since every element of $\omega$’s trajectory is in $\Omega$, $x$ is eventually iterated into $\Omega$, which makes $x$ into a pre-periodic point of $T$.
So, suppose that $V$ does not contain a cycle, and, again, let $x\in V$ be arbitrary. Since $T^{-1}\left(V\right)=V$, applying $T$ on both sides yields:
(4.5) $T\left(V\right)=T\left(T^{-1}\left(V\right)\right)\subseteq V$
So, everything in $x$’s trajectory will also be in $V$. Thus, if $x$ was pre-periodic, its trajectory would contain a cycle $\Omega$, which would contradict the assumption that $V$ contains no cycles. In particular, note that $x$’s trajectory contains a cycle if and only if there is some $y\in Y$ so that $T^{\circ n}\left(x\right)=T^{\circ m}\left(x\right)$ for two distinct positive integers $m$ and $n$; that is, cycles happen when our trajectories arrive at a point they have already visited. Thus, writing $x_{n}$ to denote $T^{\circ n}\left(x\right)$, it must be that $x_{n}\neq x_{m}$ for all distinct non-negative integers $m,n$.
So, by the pigeonhole principle, given any $R>0$, there must exist an integer $n_{R}\geq 0$ so that $\left|x_{n}\right|_{X}>R$ for all $n\geq n_{R}$. If not, then there would be infinitely many elements of $x$’s trajectory which lay in the set $\left\{y\in Y:\left|y\right|_{X}\leq R\right\}$. Since $Y$ is discrete, there are only finitely many elements in this set, which means that the trajectory of $x$ must retrace its steps at some point: there are distinct $m,n$ for which $x_{m}=x_{n}$, which, as we saw, is impossible. Since $R$ was arbitrary, we conclude that $\lim_{n\rightarrow\infty}\left|x_{n}\right|_{X}=\infty$, which proves that $x$ is a divergent point, and makes $V$ into a divergent trajectory of $T$. Since $x$ was arbitrary, we conclude that every element of $V$ is a divergent point of $T$.
Q.E.D.
5. Hydra Maps
It is a standard result of algebraic number theory that one can choose a finite set $\left\{b_{1},\ldots,b_{N}\right\}$ in $\mathcal{O}_{K}$ so that every element of $\mathcal{O}_{K}$ can be uniquely written as:
(5.1) $m_{1}b_{1}+\ldots+m_{N}b_{N}$
for some $m_{1},\ldots,m_{N}\in\mathbb{Z}$. That is, as an abelian group under addition, $\mathcal{O}_{K}$ is a lattice. In a similar way, every non-zero ideal of $\mathcal{O}_{K}$ can be realized as lattice, as well.
We use the term affine lattice to refer to a coset of a lattice in $\mathcal{O}_{K}$; thus, every affine lattice in $\mathcal{O}_{K}$ is of the form:
(5.2) $c+\Lambda\overset{\textrm{def}}{=}\left\{c+z:z\in\Lambda\right\}$
for some ideal $\Lambda\subseteq\mathcal{O}_{K}$. We call $\Lambda$ the underlying ideal of $c+\Lambda$. Note that the underlying ideal is necessarily unique, and is given by $\Lambda=\left\{z-w:z,w\in c+\Lambda\right\}$.
In this context, letting $\Lambda$ be a lattice in $\mathcal{O}_{K}$ with $p$ distinct cosets, a $p$-Hydra map $H:\mathcal{O}_{K}\rightarrow\mathcal{O}_{K}$ is given by a collection of $p$ affine linear maps, called branches, which we associate to each of the cosets of $\Lambda$. In particular, given $z\in\mathcal{O}_{K}$, $H\left(z\right)$ is obtained by applying to $z$ the branch specified by the coset of $\Lambda$ that $z$ happens to lie in.
Example 1.
The lattice in question is the ideal $2\mathbb{Z}$, and the parity of $n$ is precisely the coset of $2\mathbb{Z}$ that $n$ happens to lie in.
One difficulty that arises in generalizing maps of this shape to $\mathcal{O}_{K}$ is that while every map of the form:
(5.3) $z\in K\mapsto az+b\in K$
for fixed constants $a,b\in K$ is affine linear, not every affine linear map on $K$ will be of this form, as there exist linear maps $K\rightarrow K$ which are not given by multiplication against some fixed scalar. To deal with this, we introduce the notion of an affine lattice morphism.
Definition 8.
We write $\textrm{Aff}_{\mathbb{F}}K$ to denote the group of all invertible affine $\mathbb{F}$-linear maps $K\rightarrow K$; these are maps of the form:
(5.4) $z\mapsto rz+c$
where $r\in\textrm{End}_{\mathbb{F}}^{\times}\left(K\right)$ and $c\in K$; $\textrm{Aff}_{\mathbb{F}}K$ is a non-abelian group under map composition. We write $\left(r,c\right)$ as a shorthand for a map of the form (5.4).
Remark 3.
It is important to distinguish between $\mathbb{Q}$-linear maps on $K$ and $K$-linear maps on $K$. For the case $K=\mathbb{Q}\left(\sqrt{2}\right)$, letting $a+b\sqrt{2}\in K$ be arbitrary (where $a,b\in\mathbb{Q}$), consider the maps $m_{1},m_{2}:K\rightarrow K$ given by:
(5.5) $m_{1}\left(a+b\sqrt{2}\right)\overset{\textrm{def}}{=}$ $\sqrt{2}\left(a+b\sqrt{2}\right)$
(5.6) $m_{2}\left(a+b\sqrt{2}\right)\overset{\textrm{def}}{=}$ $a+b$
Both $m_{1}$ and $m_{2}$ are $\mathbb{Q}$-linear, but only $m_{1}$ is $K$-linear. Indeed:
(5.7) $m_{2}\left(\sqrt{2}\left(a+b\sqrt{2}\right)\right)$ $=m_{2}\left(2b+a\sqrt{2}\right)=a+2b$
(5.8) $\sqrt{2}m_{2}\left(a+b\sqrt{2}\right)$ $=\sqrt{2}\left(a+b\right)\neq a+2b$
Now we can give the definition of a Hydra map.
Definition 9.
A pre-Hydra consists of the following data:
I. A global field $K$.
II. A proper, non-zero ideal $\Lambda\subset\mathcal{O}_{K}$.
II. For each element $j\in\mathcal{O}_{K}/\Lambda$, an affine linear map $\left(r_{j},c_{j}\right)\in\textrm{Aff}_{\mathbb{F}}K$. We often denote $\left(r_{j},c_{j}\right)$ as $H_{j}$, and call $H_{j}$ the $j$th branch of the pre-Hydra.
Definition 10.
A $\Lambda$-Hydra (map) on $K$ is a map $H:\mathcal{O}_{K}\rightarrow\mathcal{O}_{K}$ defined by:
(5.9) $H\left(z\right)\overset{\textrm{def}}{=}\sum_{j\in\mathcal{O}_{K}/\Lambda}\left[z\overset{\Lambda}{\equiv}j\right]\left(r_{j}z+c_{j}\right)$
Here, $\left[z\overset{\Lambda}{\equiv}j\right]$ is the function which is $1$ if $z$ is congruent to $j$ mod $\Lambda$, and is 0 otherwise, and $\left(K,\Lambda,\left\{\left(r_{j},c_{j}\right)\right\}_{j\in\mathcal{O}_{K}/\Lambda}\right)$ is a pre-Hydra so that, for all $j\in\mathcal{O}_{K}/\Lambda$, we have $r_{j}z+c_{j}\in\mathcal{O}_{K}$ for all $z\in\mathcal{O}_{K}$.
We call $\Lambda$ the underlying ideal of $H$, and write $H_{j}$ to denote the affine linear map $\left(r_{j},c_{j}\right)$, which we call the $j$th branch of $H$.
Remark 4.
As defined, $H\left(z\right)$ applies $r_{j}z+c_{j}$, where $j$ is the representative of the coset of $\Lambda$ to which $z$ belongs. (5.9) can be written in a single line as:
(5.10) $H\left(z\right)=r_{\left[z\right]_{\Lambda}}z+c_{\left[z\right]_{\Lambda}}$
where $\left[z\right]_{\Lambda}$ denotes the image of $z$ under the canonical projection $\mathcal{O}_{K}\rightarrow\mathcal{O}_{K}/\Lambda$.
Remark 5.
As $\Lambda$ always has at least two cosets in $\mathcal{O}_{K}$, we say a map $H:\mathcal{O}_{K}\rightarrow\mathcal{O}_{K}$ is a $p$-Hydra map whenever $H$ is a $\Lambda$-Hydra map for some $\Lambda\subset\mathcal{O}_{K}$ so that $\left|\mathcal{O}_{K}/\Lambda\right|=p$.
Some notable properties of $\Lambda$-Hydras are as follows:
Definition 11.
Let $H$ be a $\Lambda$-Hydra on $K$. We say $H$ is:
I. Integral, if for all $z\in\mathcal{O}_{K}$ and all $j\in\mathcal{O}_{K}/\Lambda$, $H_{j}\left(z\right)\in\mathcal{O}_{K}$ if and only if $z\in j+\Lambda$.
II. Proper, if the map $1-r_{0}\in\textrm{End}_{\mathbb{F}}\left(K\right)$ is invertible.
III. Centered, if $H_{0}\left(0\right)=0$.
Remark 6.
From the perspective of dynamical systems, the integrality property of a Hydra is, by far, its most important property, and plays a critical role in the proof of the Correspondence Principle (Theorem 1, given at the end of Section 6).
The classic Collatz map $C:\mathbb{Z}\rightarrow\mathbb{Z}$ defined by:
$C\left(n\right)\overset{\textrm{def}}{=}\begin{cases}\frac{n}{2}&\textrm{if }n\overset{2}{\equiv}0\\ 3n+1&\textrm{if }n\overset{2}{\equiv}1\end{cases}$
is a $2$-Hydra on $\mathbb{Z}$ which is non-integral, as the “odd” branch $H_{1}\left(x\right)=3x+1$ sends both even and odd integers to integers. However, this is not a significant obstruction; the classic “shortening” of $C$ given by the aptly-named Shortened Collatz Map $T_{3}:\mathbb{Z}\rightarrow\mathbb{Z}$:
(5.11) $T_{3}\left(n\right)\overset{\textrm{def}}{=}\begin{cases}\frac{n}{2}&\textrm{if }n\textrm{ is even}\\ \frac{3n+1}{2}&\textrm{if }n\textrm{ is odd}\end{cases}$
is integral, as its “even branch” $H_{0}$ is integer-valued only for even integers, while its “odd” branch $H_{1}$ is integer-valued only for odd integers.
Remark 7.
Letting $a\in\mathcal{O}_{K}$ be undetermined, one can often conjugate an uncentered Hydra $H$ to a centered Hydra $H_{\textrm{center}}$ by writing:
$H_{\textrm{center}}\left(z\right)=H\left(z+a\right)-a$
for a judicious choice of $a\in\mathcal{O}_{K}$.
Remark 8.
The centeredness plays a role in the Correspondence Principle.
6. The Numen of a Hydra Map
Throughout this section, let $H:\mathcal{O}_{K}\rightarrow\mathcal{O}_{K}$ be a proper $\Lambda$-Hydra map. Following [12, 14], we construct $X_{H}$, the numen of $H$, like so.
Definition 12.
Let $\Sigma_{\Lambda}^{*}$ denote the set of strings (finite sequences) of elements of $\mathcal{O}_{K}/\Lambda$ including the empty string, represented by the empty set $\varnothing$. given any string $\mathbf{j}=\left(j_{1},\ldots,j_{N}\right)\in\Sigma_{\Lambda}^{*}$, we call $N$ the length of $\mathbf{j}$, and denote it by $\left|\mathbf{j}\right|$. The empty string is defined as the unique string of length 0. We write $\wedge:\Sigma_{\Lambda}^{*}\times\Sigma_{\Lambda}^{*}\rightarrow\Sigma_{\Lambda}^{*}$ to denote the concatenation operator:
(6.1) $\mathbf{i}\wedge\mathbf{j}\overset{\textrm{def}}{=}\left(i_{1},\ldots,i_{\left|\mathbf{i}\right|},j_{1},\ldots,j_{\left|\mathbf{j}\right|}\right)$
We write $\Sigma_{\Lambda}^{\infty}$ to denote the set of strings of finite or infinite length. Note that we can extend $\wedge$ to a map $\wedge:\Sigma_{\Lambda}^{*}\times\Sigma_{\Lambda}^{\infty}\rightarrow\Sigma_{\Lambda}^{\infty}$.
Definition 13.
Given any string $\mathbf{j}\in\Sigma_{\Lambda}^{*}$, we define the composition sequence $H_{\mathbf{j}}\in\textrm{Aff}K$ by:
(6.2) $H_{\mathbf{j}}\left(z\right)\overset{\textrm{def}}{=}\left(H_{j_{1}}\circ\cdots\circ H_{j_{\left|\mathbf{j}\right|}}\right)\left(z\right),\textrm{ }\forall z\in K$
When $\mathbf{j}$ is the empty string, we define $H_{\varnothing}$ to be the identity map on $K$.
As an element of $\textrm{Aff}K$, the map $H_{\mathbf{j}}$ can be written in the form $\left(r,c\right)$ for some $r\in\textrm{End}_{\mathbb{F}}^{\times}\left(K\right)$ and $c\in K$ depending on $\mathbf{j}$. We denote these by $M_{H}\left(\mathbf{j}\right)$ and $X_{H}\left(\mathbf{j}\right)$, respectively, so that:
(6.3) $H_{\mathbf{j}}\left(z\right)=M_{H}\left(\mathbf{j}\right)z+X_{H}\left(\mathbf{j}\right),\textrm{ }\forall z\in K$
In this way, we can realize $M_{H}$ and $X_{H}$ as maps $\Sigma_{\Lambda}^{*}\rightarrow\textrm{End}_{\mathbb{F}}^{\times}\left(K\right)$ and $\Sigma_{\Lambda}^{*}\rightarrow K$, respectively. $X_{H}$ is called the numen of $H$, and is defined by:
(6.4) $X_{H}\left(\mathbf{j}\right)\overset{\textrm{def}}{=}H_{\mathbf{j}}\left(0\right),\textrm{ }\forall\mathbf{j}\in\Sigma_{\Lambda}^{*}$
The next two identities are trivial, and follow immediately from the definitions:
Proposition 7.
For any $\mathbf{i},\mathbf{j}\in\Sigma_{\Lambda}^{*}$:
(6.5) $M_{H}\left(\mathbf{i}\wedge\mathbf{j}\right)=M_{H}\left(\mathbf{i}\right)M_{H}\left(\mathbf{j}\right)$
(6.6) $X_{H}\left(\mathbf{i}\wedge\mathbf{j}\right)=H_{\mathbf{i}}\left(X_{H}\left(\mathbf{j}\right)\right)$
Definition 14.
Letting $p$ denote $\left|\mathcal{O}_{K}/\Lambda\right|$, we fix once and for all choice of an enumeration $\Lambda_{0},\ldots,\Lambda_{p-1}$ of the cosets of $\Lambda$ in $\mathcal{O}_{K}$, with $\Lambda_{0}$ denoting $\Lambda$ itself and $\Lambda_{1},\ldots,\Lambda_{p-1}$ denoting its non-trivial cosets. We then let $\beta:\mathcal{O}_{K}/\Lambda\rightarrow\left\{0,\ldots,p-1\right\}$ be the map which sends $i\in\mathcal{O}_{K}/\Lambda$ to the unique integer $\beta\left(i\right)\in\left\{0,\ldots,p-1\right\}$ so that $i\in\Lambda_{\beta\left(i\right)}$. (Thus, if $i\in\Lambda_{0}$, then $\beta\left(i\right)=0$; if $i\in\Lambda_{1}$, then $\beta\left(i\right)=1$, etc.).
Because of this, we shall write $\Sigma_{p}^{*}$ instead of $\Sigma_{\Lambda}^{*}$, with $\Sigma_{p}^{*}$ denoting the set of all finite strings of the numbers $\left\{0,\ldots,p-1\right\}$, including the empty string. $\beta$ then induces a bijection between $\Sigma_{\Lambda}^{*}$ and $\Sigma_{p}^{*}$.
Letting $\mathbb{N}_{0}=\left\{0,1,2,3,\ldots\right\}$ denote the set of non-negative integers, we then write $\textrm{DigSum}_{p}:\Sigma_{p}^{*}\rightarrow\mathbb{N}_{0}$ to denote the function:
(6.7) $\textrm{DigSum}_{p}\left(\mathbf{j}\right)=\sum_{n=1}^{\left|\mathbf{j}\right|}\beta\left(j_{n}\right)p^{n-1}$
where $\left|\mathbf{j}\right|$ is the length of $\mathbf{j}$ (the number of entries of $\mathbf{j}$).
Remark 9.
Like with writing $\Sigma_{p}^{*}$ instead of $\Sigma_{\Lambda}^{*}$, going forward, in an very important abuse of notation, we will index indexing $r_{j}$ and $c_{j}$ using $j\in\left\{0,\ldots,p-1\right\}$ by way of $\beta$, instead of using $j\in\mathcal{O}_{K}/\Lambda$. Thus, for example, by $r_{1}$, we actually mean $r_{\beta^{-1}\left(1\right)}$, where $\beta^{-1}\left(1\right)$ is the unique $j\in\mathcal{O}_{K}/\Lambda$ that $\beta$ sends to $1$.
Example 2.
If $\mathcal{O}_{K}=\mathbb{Z}$ and $\Lambda=3\mathbb{Z}$, then upon identifying $\mathcal{O}_{K}/\Lambda=\mathbb{Z}/3\mathbb{Z}$ with $\left\{0,1,2\right\}$, we have that:
(6.8) $\textrm{DigSum}_{3}\left(\mathbf{j}\right)=\sum_{n=1}^{\left|\mathbf{j}\right|}j_{n}3^{n-1}$
Thus, for example:
(6.9) $\textrm{DigSum}_{3}\left(\left(1,2,0,1\right)\right)=1+2\cdot 3+0\cdot 3^{2}+1\cdot 3^{3}$
Although $\textrm{DigSum}_{p}:\Sigma_{p}^{*}\rightarrow\mathbb{N}_{0}$ is surjective, note that it is not injective. Indeed, given any $\mathbf{i}\in\Sigma_{p}^{*}$ if we append finitely many 0s to the right of $\mathbf{i}$, we obtain a string $\mathbf{j}\neq\mathbf{i}$ for which $\textrm{DigSum}_{p}\left(\mathbf{i}\right)=\textrm{DigSum}_{p}\left(\mathbf{j}\right)$. To that end:
Definition 15.
Let $\sim$ be the equivalence relation on $\Sigma_{p}^{*}$ defined by $\mathbf{i}\sim\mathbf{j}$ if and only if $\textrm{DigSum}_{p}\left(\mathbf{i}\right)=\textrm{DigSum}_{p}\left(\mathbf{j}\right)$. We then write $\Sigma_{p}^{*}/\sim$ to denote the set of equivalence classes in $\Sigma_{p}^{*}$ under $\sim$.
This gives the easy result:
Proposition 8.
The map $\Sigma_{p}^{*}/\sim\rightarrow\mathbb{N}_{0}$ induced by $\textrm{DigSum}_{p}$ is a bijection.
With this, we can realize $X_{H}$ as a function $\mathbb{N}_{0}\rightarrow K$.
Proposition 9.
$X_{H}\left(\mathbf{i}\right)=X_{H}\left(\mathbf{j}\right)$ for all $\mathbf{i}$ and $\mathbf{j}$ belonging to the same equivalence class in $\Sigma_{p}^{*}/\sim$. Hence, $X_{H}:\Sigma_{p}^{*}/\sim\rightarrow K$ is well-defined. We can then view $X_{H}$ as a function $\mathbb{N}_{0}\rightarrow K$ by using $\textrm{DigSum}_{p}$ to identify $\Sigma_{p}^{*}/\sim$ with $\mathbb{N}_{0}$.
Using the definition of $X_{H}\left(\mathbf{j}\right)$ as $H_{\mathbf{j}}\left(0\right)$, one can easily give the following functional equation characterization of $X_{H}:\mathbb{N}_{0}\rightarrow K$.
Proposition 10 (Siegel (2022); see [12]).
$X_{H}:\mathbb{N}_{0}\rightarrow K$ is the unique function satisfying the functional equations:
(6.10) $X_{H}\left(pn+j\right)=r_{j}X_{H}\left(n\right)+c_{j},\textrm{ }\forall n\in\mathbb{N}_{0},\textrm{ }\forall j\in\left\{0,\ldots,p-1\right\}$
where $r_{j}X_{H}\left(n\right)$ is the image of $X_{H}\left(n\right)\in K$ under $r_{j}\in\textrm{End}_{\mathbb{F}}^{\times}\left(K\right)$.
Remark 10.
Proposition 10 only requires $H$ to be proper. Integrality and centeredness are not needed.
Next, we introduce the ring of $p$-adic integers $\mathbb{Z}_{p}$.
Definition 16.
Given an integer $p\geq 2$, the ring of $p$-adic integers is defined by the following quotient of a formal power series ring:
(6.11) $\mathbb{Z}_{p}\overset{\textrm{def}}{=}\frac{\mathbb{Z}\left[\left[x\right]\right]}{\left\langle x-p\right\rangle}$
In practice, one can think of a $p$-adic integer as an integer written in base $p$, but with the caveat that said integer can potentially have infinitely many base $p$ digits. The quotient construction justifies viewing $p$-adic integers as “formal power series in $p$ with coefficients in $\left\{0,\ldots,p-1\right\}$”. Given $\mathfrak{z}\in\mathbb{Z}_{p}$, this power series representation of $\mathfrak{z}$ is called the Hensel series of $\mathfrak{z}$, and the coefficients of this series are called the $p$-adic digits of $\mathfrak{z}$. $\mathbb{Z}_{p}$ is made into a compact abelian group by equipping it with the adic topology induced by the maximal ideal xpx-p of $\mathbb{Z}\left[\left[x\right]\right]$.
When $p$ is a prime number, $\mathbb{Z}_{p}$ can be topologized using the $p$-adic absolute value. This is obtained by defining the $p$-adic valuation $v_{p}:\mathbb{Q}\rightarrow\mathbb{Z}\cup\left\{+\infty\right\}$ by:
(6.12) $v_{p}\left(n\right)=\#\textrm{ of times }p\textrm{ divides }n$
for all positive integers $n$, and then using the identities:
(6.13) $v_{p}\left(\frac{1}{n}\right)=-v_{p}\left(n\right),\textrm{ }\forall n\geq 1$
and
(6.14) $v_{p}\left(rs\right)=v_{p}\left(r\right)+v_{p}\left(s\right),\textrm{ }\forall r,s\in\mathbb{Q}$
to extend $v_{p}$ to $\mathbb{Q}$. Note that $v_{p}\left(0\right)$ is defined to be $\infty$. The $p$-adic absolute value is then defined by:
(6.15) $\left|r\right|_{p}\overset{\textrm{def}}{=}p^{-v_{p}\left(r\right)},\textrm{ }\forall r\in\mathbb{Q}$
In this way, $\mathbb{Z}_{p}$ can be constructed as the completion of the metric space obtained by equipping $\mathbb{Z}$ with the metric induced by the $p$-adic absolute value. $\mathbb{Q}_{p}$, the field of $p$-adic numbers, is the field of fractions of $\mathbb{Z}_{p}$, or, equivalently, the completion of $\mathbb{Q}$ with respect to the metric induced by $\left|\cdot\right|_{p}$. Importantly, the $p$-adic absolute value is non-archimedean, meaning it satisfies the ultrametric inequality:
(6.16) $\left|\mathfrak{x}+\mathfrak{y}\right|_{p}\leq\max\left\{\left|\mathfrak{x}\right|_{p},\left|\mathfrak{y}\right|_{p}\right\},\textrm{ }\forall\mathfrak{x},\mathfrak{y}\in\mathbb{Q}_{p}$
with equality whenever $\left|\mathfrak{x}\right|_{p}\neq\left|\mathfrak{y}\right|_{p}$.
Finally, given any $\mathfrak{z}\in\mathbb{Z}_{p}$ and any integer $n\geq 0$, we write $\left[\mathfrak{z}\right]_{p^{n}}$ to denote the unique integer in $\left\{0,\ldots,p^{n}-1\right\}$ which is congruent to $\mathfrak{z}$ mod $p^{n}$.
Remark 11.
Using this definition, it can be shown that, as a compact abelian group, $\mathbb{Z}_{p}$ is isomorphic to the direct product:
(6.17) $\prod_{\ell\mid p}\mathbb{Z}_{\ell}$
taken over all prime divisors $\ell$ of $p$. In particular, this shows that for any prime $\ell$ and any integer $\geq 1, n\mathbb{Z}_{\ell^{n}}$ is isomorphic to $\mathbb{Z}_{\ell}$ as a topological ring. As such, we will write the above direct product isomorphism as:
(6.18) $\mathbb{Z}_{p}\cong\prod_{\ell\mid p}\mathbb{Z}_{\ell^{v_{\ell}\left(p\right)}}$
Ex:
(6.19) $\mathbb{Z}_{12}\cong\mathbb{Z}_{3}\times\mathbb{Z}_{4}$
Just like with $\mathbb{N}_{0}$ and $\Sigma_{\Lambda}^{*}$, we can identify $\mathbb{Z}_{p}$ with a set of strings.
Definition 17.
We write:
I. $\Sigma_{p}^{\infty}$, to denote the set of all strings of elements of $\left\{0,\ldots,p-1\right\}$ of either finite or infinite length. We also include the empty string as an element of $\Sigma_{p}^{\infty}$.
II. $\Sigma_{p}^{+\infty}$, to denote the set of all strings of elements of $\left\{0,\ldots,p-1\right\}$ of infinite length.
Remark 12.
We can and will extend $\textrm{DigSum}_{p}$ from a function $\Sigma_{p}^{*}\rightarrow\mathbb{N}_{0}$ to a function $\Sigma_{p}^{\infty}\rightarrow\mathbb{Z}_{p}$, where $\Sigma_{p}^{\infty}$ is the set of finite- or infinite-length strings in the symbols $\left\{0,\ldots,p-1\right\}$. Letting $\sim$ denote the equivalence relation on $\Sigma_{p}^{\infty}$ with $\mathbf{i}\sim\mathbf{j}$ if and only if $\textrm{DigSum}_{p}\left(\mathbf{i}\right)=\textrm{DigSum}_{p}\left(\mathbf{j}\right)$, we then have that $\textrm{DigSum}_{p}$ induces a bijection $\Sigma_{p}^{\infty}/\sim\rightarrow\mathbb{Z}_{p}$, where $\Sigma_{p}^{\infty}/\sim$ is the set of equivalence classes of $\Sigma_{p}^{\infty}$ under $\sim$.
In [12], it was shown that, given mild assumptions on the $r_{j}$s, one can use the identification of $\Sigma_{p}^{\infty}/\sim$ with $\mathbb{Z}_{p}$ to interpolate $X_{H}:\mathbb{N}_{0}\rightarrow K$ to a function on $\mathbb{Z}_{p}$. The details for this are given in Lemma 1, below. The essential idea is that by making a judicious choice of absolute values on $K$, one can realize $X_{H}$ as a function (or, in general, as a distribution) out of $\mathbb{Z}_{p}$ whose regularity depends on how close the $r_{j}$s are to behaving as contractions with respect to $K$. The behavior of the $r_{j}$s is captured by their norm with respect to a given absolute value of $K$. The result given below is a slightly modified version of the versions originally stated in [14]. The reader should be aware that Lemma 1 treats $\mathbb{Z}_{p}$ as a measure space equipped with its real-valued Haar probability measure, and treat $K_{\ell}$ as a measure space equipped with its real-valued Haar measure. We also use the concept of a locally constant function. Readers unfamiliar with either of these can find an expository treatment given of them at the start of Section 7.
Lemma 1.
For each non-trivial place $\ell$ of $K$, let:
(6.20) $\rho_{H,\ell}\overset{\textrm{def}}{=}\prod_{j=0}^{p-1}\left\|r_{j}\right\|_{\ell}$
be the product of the $\ell$-adic norms of the $r_{j}$s.
Now, consider the extension $X_{H}:\mathbb{Z}_{p}\rightarrow K_{\ell}$ defined by the formula
(6.21) $X_{H}\left(\mathfrak{z}\right)$ $\overset{K_{\ell}}{=}\lim_{n\rightarrow\infty}X_{H}\left(\left[\mathfrak{z}\right]_{p^{n}}\right)$
for all $\mathfrak{z}\in\mathbb{Z}_{p}$ for which $X_{H}\left(\mathfrak{z}\right)$ exists. Then:
I. If $\rho_{H,\ell}<1$, then (6.21) defines a measurable function $X_{H}:\mathbb{Z}_{p}\rightarrow K_{\ell}$, with (6.21) converging for almost every $\mathfrak{z}\in\mathbb{Z}_{p}$. In particular, it converges at all $\mathfrak{z}\in\mathbb{Z}_{p}$ for which:
(6.22) $\sum_{j=0}^{p-1}d_{H,\ell,j}\left(\mathfrak{z}\right)\ln\left\|r_{j}\right\|_{\ell}<0$
where:
(6.23) $d_{H,\ell,j}\left(\mathfrak{z}\right)\overset{\textrm{def}}{=}\begin{cases}\limsup_{m\rightarrow\infty}\frac{\#_{j}\left(\left[\mathfrak{z}\right]_{p^{m}}\right)}{m}&\textrm{if }\left\|r_{j}\right\|_{\ell}\leq 1\\ \liminf_{m\rightarrow\infty}\frac{\#_{j}\left(\left[\mathfrak{z}\right]_{p^{m}}\right)}{m}&\textrm{if }\left\|r_{j}\right\|_{\ell}>1\end{cases}$
is the $\left(H,\ell\right)$-normalized upper density of $j$ in the $p$-adic digits of $\mathfrak{z}$.
II. If $\rho_{H,\ell}<1$ and $\max_{0\leq j<p}\left\|r_{j}\right\|_{\ell}<1$, then (6.21) converges uniformly with respect to $\mathfrak{z}\in\mathbb{Z}_{p}$, making $X_{H}:\mathbb{Z}_{p}\rightarrow K_{\ell}$ into a continuous function.
Moreover, given either (I) or (II), we have that:
(6.24) $X_{H}\left(p\mathfrak{z}+j\right)=r_{j}X_{H}\left(\mathfrak{z}\right)+c_{j}$
occurs for all $j\in\left\{0,\ldots,p-1\right\}$ and all $\mathfrak{z}\in\mathbb{Z}_{p}$ for which $X_{H}\left(\mathfrak{z}\right)$ converges in $K_{\ell}$, and that the extensions of $X_{H}$ given in (I) and (II), respectively, are the unique functions satisfying these functional equations. Furthermore, if $\ell$ is non-archimedean and $\max_{0\leq j<p}\left\|r_{j}\right\|_{\ell}\leq 1$, then:
(6.25) $\limsup_{N\rightarrow\infty}\sup_{\mathfrak{z}\in\mathbb{Z}_{p}}\left|X_{H}\left(\left[\mathfrak{z}\right]_{p^{N}}\right)\right|_{\ell}\leq\max_{0\leq j<p}\left|c_{j}\right|_{\ell}$
In particular, $X_{H}:\mathbb{Z}_{p}\rightarrow K_{\ell}$ is $\ell$-adically bounded by $\max_{0\leq j<p}\left|c_{j}\right|_{\ell}$ on its set of convergence.
Remark 13.
Like with Proposition 10, Lemma 1 requires only the properness of $H$, as is made clear near the end of the proof.
Proof: First, note that if $\mathfrak{z}\in\mathbb{N}_{0}$, then $\left[\mathfrak{z}\right]_{p^{n}}=\mathfrak{z}$ for all sufficiently large $n$, and hence (6.21) converges. So, suppose $\mathfrak{z}\in\mathbb{Z}_{p}\backslash\mathbb{N}_{0}$. Next, following [14], a simple inductive argument gives the following formula for the value of $X_{H}$ at a given string $\mathbf{j}$ of finite length:
(6.26) $X_{H}\left(\mathbf{j}\right)=\sum_{m=0}^{\left|\mathbf{j}\right|-1}r_{\left[\mathbf{j}\right]_{m}}c_{j_{m+1}},\textrm{ }\forall\mathbf{j}\in\Sigma_{p}^{*}$
where:
(6.27) $r_{\left[\mathbf{j}\right]_{m}}\overset{\textrm{def}}{=}\prod_{k=1}^{m}r_{j_{k}}=\begin{cases}\textrm{Id}&\textrm{if }\mathbf{j}=\varnothing\\ r_{j_{1}}\cdots r_{j_{m}}&\textrm{if }\left|\mathbf{j}\right|\geq 1\end{cases}$
Applying norms, and using
(6.28) $\left|r_{j}z\right|_{\ell}\leq\left\|r_{j}\right\|_{\ell}\left|z\right|_{\ell},\textrm{ }\forall z\in K,\textrm{ }\forall j\in\left\{0,\ldots,p-1\right\}$
the triangle inequality gives:
(6.29) $\left|X_{H}\left(\mathbf{j}\right)\right|_{\ell}=\sum_{m=0}^{\left|\mathbf{j}\right|-1}\left|c_{j_{m+1}}\right|_{\ell}\prod_{k=1}^{m}\left\|r_{j_{k}}\right\|_{\ell}$
So, given $\mathfrak{z}\in\mathbb{Z}_{p}$, for each $n\geq 0$, let:
(6.30) $\mathbf{j}_{n}\left(\mathfrak{z}\right)=\left(j_{1}\left(\mathfrak{z}\right),j_{2}\left(\mathfrak{z}\right),\ldots,j_{n}\left(\mathfrak{z}\right)\right)$
denote the string of $p$-adic digits of $\left[\mathfrak{z}\right]_{p^{n}}$; that is:
(6.31) $\left[\mathfrak{z}\right]_{p^{n}}=\textrm{DigSum}_{p}\left(\mathbf{j}_{n}\left(\mathfrak{z}\right)\right)$
Thus, we have:
(6.32) $\left|X_{H}\left(\left[\mathfrak{z}\right]_{p^{n}}\right)\right|_{\ell}=\left|X_{H}\left(\mathbf{j}_{n}\left(\mathfrak{z}\right)\right)\right|_{\ell}\ll\sum_{m=0}^{n-1}\prod_{k=1}^{m}\left\|r_{j_{k}\left(\mathfrak{z}\right)}\right\|_{\ell}=\sum_{m=0}^{n-1}\prod_{j=0}^{p-1}\left\|r_{j}\right\|_{\ell}^{\#_{j}\left(\left[\mathfrak{z}\right]_{p^{m}}\right)}$
where $_{j}\left(\left[\mathfrak{z}\right]_{p^{m}}\right)$ counts the number of digits of $\left[\mathfrak{z}\right]_{p^{m}}$ which are equal to $j$. Hence:
(6.33) $\left|X_{H}\left(\left[\mathfrak{z}\right]_{p^{n}}\right)\right|_{\ell}\ll\sum_{m=0}^{n-1}\prod_{j=0}^{p-1}\left\|r_{j}\right\|_{\ell}^{\#_{j}\left(\left[\mathfrak{z}\right]_{p^{m}}\right)}$
where the constant of proportionality is $C$.
Applying the Root Test for series convergence, and noting that $0\leq\rho_{j}\leq 1$ for all $j$, we see that the upper bound in (6.33) will converge provided:
(6.34) $\limsup_{m\rightarrow\infty}\prod_{j=0}^{p-1}\left\|r_{j}\right\|_{\ell}^{\frac{\#_{j}\left(\left[\mathfrak{z}\right]_{p^{m}}\right)}{m}}<1$
Since, for almost every $\mathfrak{z}\in\mathbb{Z}_{p}$ (in the sense of $\mathbb{Z}_{p}$’s Haar probability measure), one has:
(6.35) $\lim_{m\rightarrow\infty}\frac{\#_{j}\left(\left[\mathfrak{z}\right]_{p^{m}}\right)}{m}=\frac{1}{p}$
for all $j\in\left\{0,\ldots,p-1\right\}$, (6.34) is equivalent to:
(6.36) $\prod_{j=0}^{p-1}\left\|r_{j}\right\|_{\ell}^{1/p}<1$
The left-hand side is $\rho_{H,\ell}^{1/p}$. Thus, $\rho_{H,\ell}<1$ implies almost-everywhere convergence. As $X_{H}$ is defined by the limit $\lim_{n\rightarrow\infty}X_{H}\left(\left[\mathfrak{z}\right]_{p^{n}}\right)$, and since $X_{H}\left(\left[\mathfrak{z}\right]_{p^{n}}\right)$ is a locally constant, and hence, measurable $K_{\ell}$-valued function on $\mathbb{Z}_{p}$, it follows that the almost-everywhere convergent limit:
(6.37) $X_{H}\left(\mathfrak{z}\right)\overset{K_{\ell}}{=}\lim_{n\rightarrow\infty}X_{H}\left(\left[\mathfrak{z}\right]_{p^{n}}\right)$
is measurable, as well.
Finally, if $\max_{0\leq j<p}\left\|r_{j}\right\|_{\ell}<1$, set:
(6.38) $\rho=\max_{0\leq j<p}\left\|r_{j}\right\|_{\ell}$
Then, we re-write (6.26) like so:
(6.39) $X_{H}\left(\left[\mathfrak{z}\right]_{p^{n}}\right)=\sum_{m=0}^{n-1}\left(\prod_{k=0}^{m-1}r_{\left[\theta_{p}^{\circ k}\left(\mathfrak{z}\right)\right]_{p}}\right)c_{\left[\theta_{p}^{\circ m}\left(\mathfrak{z}\right)\right]_{p}},\textrm{ }\forall\mathfrak{z}\in\mathbb{Z}_{p},\textrm{ }\forall n\geq 0$
Here, $\theta_{p}:\mathbb{Z}_{p}\rightarrow\mathbb{Z}_{p}$ is the $p$-adic shift map, which acts on Hensel series representations as:
(6.40) $\theta_{p}\left(\sum_{n=0}^{\infty}d_{n}p^{n}\right)\overset{\textrm{def}}{=}\sum_{n=0}^{\infty}d_{n+1}p^{n}$
i.e.:
(6.41) $\theta_{p}\left(\mathfrak{z}\right)=\frac{\mathfrak{z}-\left[\mathfrak{z}\right]_{p}}{p}$
$\left[\theta_{p}^{\circ k}\left(\mathfrak{z}\right)\right]_{p}$, the projection mod $p$ of the image of $\mathfrak{z}$ under $k$ applications of $\theta_{p}$, is then precisely the coefficient of $p^{k}$ in $\mathfrak{z}$’s Hensel series representation. This then gives us the formal expression:
(6.42) $\lim_{N\rightarrow\infty}X_{H}\left(\left[\mathfrak{z}\right]_{p^{N}}\right)=\sum_{m=0}^{\infty}\left(\prod_{k=0}^{m-1}r_{\left[\theta_{p}^{\circ k}\left(\mathfrak{z}\right)\right]_{p}}\right)c_{\left[\theta_{p}^{\circ m}\left(\mathfrak{z}\right)\right]_{p}}$
We then have:
$\left|X_{H}\left(\left[\mathfrak{z}\right]_{p^{n}}\right)-\lim_{N\rightarrow\infty}X_{H}\left(\left[\mathfrak{z}\right]_{p^{N}}\right)\right|_{\ell}$ $=\left|\sum_{m=n}^{\infty}\left(\prod_{k=0}^{m-1}r_{\left[\theta_{p}^{\circ k}\left(\mathfrak{z}\right)\right]_{p}}\right)c_{\left[\theta_{p}^{\circ m}\left(\mathfrak{z}\right)\right]_{p}}\right|_{\ell}$
$\leq\sum_{m=n}^{\infty}\prod_{k=0}^{m-1}\left\|r_{\left[\theta_{p}^{\circ k}\left(\mathfrak{z}\right)\right]_{p}}\right\|_{\ell}\left|c_{\left[\theta_{p}^{\circ m}\left(\mathfrak{z}\right)\right]_{p}}\right|_{\ell}$
$\leq\sum_{m=n}^{\infty}\prod_{k=0}^{m-1}\rho\left(\max_{0\leq j<p}\left|c_{j}\right|_{\ell}\right)$
$\ll\sum_{m=n}^{\infty}\rho^{m}$
$\left(\rho<1\right);$ $\overset{\mathbb{R}}{=}\frac{\rho^{n}}{1-\rho}$
As $n\rightarrow\infty$, the upper bound tends to 0 uniformly with respect to $\mathfrak{z}\in\mathbb{Z}_{p}$, thereby establishing the uniform convergence of $\lim_{N\rightarrow\infty}X_{H}\left(\left[\mathfrak{z}\right]_{p^{N}}\right)$. As locally constant functions out of $\mathbb{Z}_{p}$ are continuous, this then proves that:
(6.43) $X_{H}\left(\mathfrak{z}\right)\overset{\textrm{def}}{=}\lim_{N\rightarrow\infty}X_{H}\left(\left[\mathfrak{z}\right]_{p^{N}}\right)$
is continuous.
Almost done, observing the identity:
(6.44) $X_{H}\left(\left[p\mathfrak{z}+j\right]_{p^{n}}\right)=X_{H}\left(p\left[\mathfrak{z}\right]_{p^{n-1}}+j\right)=r_{j}X_{H}\left(\left[\mathfrak{z}\right]_{p^{n-1}}\right)+\mathbf{c}_{j}$
for all $n\geq 1$, all $\mathfrak{z}\in\mathbb{Z}_{p}$, and all $j\in\left\{0,\ldots,p-1\right\}$, we see that the functional equations (6.24) are satisfied for all $\mathfrak{z}$ for which $\lim_{n\rightarrow\infty}X_{H}\left(\left[\mathfrak{z}\right]_{p^{n}}\right)$ exists.
For the uniqueness, given any function $f:\mathbb{N}_{0}\rightarrow K_{\ell}$ satisfying:
(6.45) $f\left(pn+j\right)=r_{j}f\left(n\right)+c_{j},\textrm{ }\forall n\geq 0,\textrm{ }\forall j\in\left\{0,\ldots,p-1\right\}$
observe that since $1-r_{0}$ is invertible by the properness of $H$, we have for $n=j=0$:
(6.46) $f\left(0\right)=r_{0}f\left(0\right)+c_{0}$
and hence:
(6.47) $f\left(0\right)=\left(1-r_{0}\right)^{-1}c_{0}$
By induction, it follows that this value of $f\left(0\right)$ then uniquely determines $f\left(n\right)$ for all $n\geq 0$. This shows that the restriction of $X_{H}$ to $\mathbb{N}_{0}$ is unique, and hence, that the continuation of $X_{H}$ to $\mathbb{Z}_{p}$ is also unique.
Finally, suppose that $\ell$ is non-archimedean, and that the maximum of the $\ell$-adic norms of the $r_{j}$s is $\leq 1$. Then (6.42) and the ultrametric property of the $\ell$-adic norm implies that:
(6.48) $\limsup_{N\rightarrow\infty}\left|X_{H}\left(\left[\mathfrak{z}\right]_{p^{N}}\right)\right|_{\ell}\leq\max_{0\leq j<p}\left|c_{j}\right|_{\ell}$
which shows that
Q.E.D.
Remark 14.
As constructed, note that for any $\mathfrak{z}\in\mathbb{Z}_{p}\cap\mathbb{Q}$ (that is, for any rational $p$-adic integer) for which $X_{H}\left(\mathfrak{z}\right)$ is finite-valued, $X_{H}\left(\mathfrak{z}\right)$ will necessarily be an element of $K$. Since $X_{H}\left(\mathfrak{z}\right)$ is the limit of $X_{H}\left(\left[\mathfrak{z}\right]_{p^{n}}\right)$ as $n\rightarrow\infty$, one can use (6.26) to write:
(6.49) $X_{H}\left(\mathfrak{z}\right)=\sum_{m=0}^{\infty}\left(\prod_{k=0}^{m-1}r_{d_{k}\left(\mathfrak{z}\right)}\right)c_{d_{m}\left(\mathfrak{z}\right)}$
where:
(6.50) $\mathfrak{z}=\sum_{n=0}^{\infty}d_{n}\left(\mathfrak{z}\right)p^{n}$
is the Hensel series of $\mathfrak{z}$, and where:
(6.51) $\prod_{k=0}^{m-1}r_{d_{k}\left(\mathfrak{z}\right)}\overset{\textrm{def}}{=}r_{d_{0}\left(\mathfrak{z}\right)}\circ\cdots\circ r_{d_{m-1}\left(\mathfrak{z}\right)}$
with the product being the identity map whenever $m=0$. Because $\mathfrak{z}$ is rational if and only if its $p$-adic digits are eventually periodic, the eventual-periodicity of the $d_{m}$s guarantees that there will be an $M$, depending on $\mathfrak{z}$, so that the series formed by summing the terms of (6.49) for $m\geq M$ is a $K$-linear combination of finitely many geometric series. If $X_{H}\left(\mathfrak{z}\right)$ is convergent, these geometric series will necessarily converge to elements of $K$.
Remark 15.
In the study of iterated function systems (IFS), one traditionally use spaces of strings—usually called words in that subject—where, following [12], we use the $p$-adic integers. In passing from arbitrary strings $\mathbf{j}\in\Sigma_{p}^{\infty}$to the $p$-adic integers using $\textrm{DigSum}_{p}$, we lose distinction between those strings that end in infinitely many 0s. Given any iterated function system generated by maps $f_{1},\ldots,f_{N}$ on a complete normed vector space $V$, we can avoid this lost information by defining $p$ to be $N+1$, and then setting $H_{0}$ to be the identity map of $V$ and setting $H_{j}=f_{j}$ for all $j\in\left\{1,\ldots,p-1\right\}$. The numen:
(6.52) $X_{H}\left(p\mathfrak{z}+j\right)=H_{j}\left(X_{H}\left(\mathfrak{z}\right)\right),\textrm{ }\forall j\in\left\{0,\ldots,p-1\right\},\textrm{ }\forall\mathfrak{z}\in\mathbb{Z}_{p}$
is then a surjective map from $\mathbb{Z}_{p}$ onto the fractal attractor $F$ of the $f_{j}$s. If we set $q=N$ and define:
(6.53) $Y_{H}\left(q\mathfrak{z}+j\right)=f_{j+1}\left(Y_{H}\left(\mathfrak{z}\right)\right),\textrm{ }\forall j\in\left\{0,\ldots,q-1\right\},\textrm{ }\forall\mathfrak{z}\in\mathbb{Z}_{q}$
the image of $Y_{H}$ would omit all the points of $F$ corresponding to composition sequences:
(6.54) $f_{j_{1}}\circ f_{j_{2}}\circ\cdots$
so that $j_{n}=1$ for all sufficiently large $n$.
Finally, we have the analogue of the Correspondence Principle from [14].
Theorem 1 (Correspondence Principle).
Let $H:\mathcal{O}_{K}\rightarrow\mathcal{O}_{K}$ be a proper, centered, integral $\Lambda$-Hydra map. Suppose:
i. For each $n\geq 1$, there is a place $\ell_{n}$ of $K$ so that $\left\|M_{H}\left(n\right)\right\|_{\ell_{n}}<1$.
ii. There is a non-trivial non-archimedean place $v$ of $K$ so that $\left\|r_{j}\right\|_{v}>1$ for all $j\in\left\{0,\ldots,p-1\right\}$ and $v\left(\Lambda\right)\geq 1$.
iii. $H_{0}\left(0\right)=0$.
Then:
I. $z\in\mathcal{O}_{K}$ is a periodic point of $H$ if and only if there is a rational $p$-adic integer $\mathfrak{z}\in\left(\mathbb{Z}_{p}\backslash\mathbb{N}_{0}\right)\cap\mathbb{Q}$ so that $X_{H}\left(\mathfrak{z}\right)=z$.
II. If $K$ is a number field and $X_{H}\left(\mathfrak{z}\right)$ is in $\mathcal{O}_{K}$ for some irrational $p$-adic integer $\mathfrak{z}\in\mathbb{Z}_{p}\backslash\mathbb{Q}$, $X_{H}\left(\mathfrak{z}\right)$ is a divergent trajectory of $H$ with respect to any non-trivial archimedean absolute value on $K$.
Proof (Sketch):
I. Let $\mathbf{j}\in\Sigma_{p}^{*}$ be a string so that $H_{\mathbf{j}}\left(z\right)=z$, and let $n=\textrm{DigSum}_{p}\left(\mathbf{j}\right)$. Then:
(6.55) $z=H_{\mathbf{j}}\left(z\right)=M_{H}\left(n\right)z+X_{H}\left(n\right)$
Consequently, for all $m\geq 1$:
(6.56) $z=H_{\mathbf{j}}^{\circ m}\left(z\right)=M_{q}\left(\mathbf{j}^{\wedge m}\right)z+X_{H}\left(\mathbf{j}^{\wedge m}\right)=\left(M_{q}\left(n\right)\right)^{m}z+X_{H}\left(\mathbf{j}^{\wedge m}\right)$
where $\mathbf{j}^{\wedge m}$ is the string obtained by concatenating $m$ copies of $\mathbf{j}$. Here:
(6.57) $X_{H}\left(\mathbf{j}^{\wedge m}\right)=H_{\mathbf{j}}\left(X_{H}\left(\mathbf{j}^{\wedge m-1}\right)\right)=M_{H}\left(\mathbf{j}\right)X_{H}\left(\mathbf{j}^{\wedge m-1}\right)+X_{H}\left(\mathbf{j}\right)$
from which it follows that:
(6.58) $X_{H}\left(\mathbf{j}^{\wedge m}\right)=\frac{1-M_{H}^{m}\left(\mathbf{j}\right)}{1-M_{H}\left(\mathbf{j}\right)}X_{H}\left(\mathbf{j}\right)=\frac{1-M_{H}^{m}\left(n\right)}{1-M_{H}\left(n\right)}X_{H}\left(n\right)$
By (i), we can choose a place $\ell_{n}$ so that $\left\|M_{H}^{m}\left(n\right)\right\|_{\ell_{n}}=\left\|M_{H}\left(n\right)\right\|_{\ell_{n}}^{m}$ tends to 0 as $m\rightarrow\infty$. This gives us:
(6.59) $z=H_{\mathbf{j}}^{\circ m}\left(z\right)=\left(M_{q}\left(n\right)\right)^{m}z+\frac{1-M_{H}^{m}\left(n\right)}{1-M_{H}\left(n\right)}X_{H}\left(n\right),\textrm{ }\forall m\geq 1$
and hence, we get $\ell$-adic convergence:
(6.60) $z=\frac{X_{H}\left(n\right)}{1-M_{H}\left(n\right)}$
as $m\rightarrow\infty$. As in [14], we have:
(6.61) $\frac{X_{H}\left(n\right)}{1-M_{H}\left(n\right)}=X_{H}\left(B_{p}\left(n\right)\right),\textrm{ }\forall n\geq 1$
where:
(6.62) $B_{p}\left(n\right)\overset{\textrm{def}}{=}\frac{n}{1-p^{\lambda_{p}\left(n\right)}}$
is the function which outputs the $p$-adic integer whose sequence of digits is precisely the result of concatenating infinitely many copies of the sequence of $n$’s $p$-adic digits. Since $B_{p}\left(n\right)\in\left(\mathbb{Z}_{p}\backslash\mathbb{N}_{0}\right)\cap\mathbb{Q}$ for all $n\geq 1$, this shows that $z$ is in the image of $\left(\mathbb{Z}_{p}\backslash\mathbb{N}_{0}\right)\cap\mathbb{Q}$ under $X_{H}$.
For the other direction, let $z=X_{H}\left(\mathfrak{z}\right)$ be an element of $\mathcal{O}_{K}$ for some $\mathfrak{z}\in\left(\mathbb{Z}_{p}\backslash\mathbb{N}_{0}\right)\cap\mathbb{Q}$. (As for convergence, note that (i) implies that $X_{H}\left(\mathfrak{z}\right)=\lim_{N\rightarrow\infty}X_{H}\left(\left[\mathfrak{z}\right]_{p^{N}}\right)$ converges in $K_{\ell_{n}}$ to $z$ for some $n\geq 0$.) Noting the identity:
(6.63) $X_{H}\left(\mathfrak{z}\right)=H_{\left[\mathfrak{z}\right]_{p}}\left(X_{H}\left(\theta_{p}\left(\mathfrak{z}\right)\right)\right)$
which follows from rewriting $X_{H}$’s functional equation (6.24), note that since $\mathfrak{z}$ is in $\left(\mathbb{Z}_{p}\backslash\mathbb{N}_{0}\right)\cap\mathbb{Q}$, there is an $n\geq 0$ so that $\theta_{p}^{\circ n}\left(\mathfrak{z}\right)$’s sequence of $p$-adic digits is periodic with period $\tau$, being generated by a string $\mathbf{j}$ of length $\tau$. As such, letting $\mathfrak{x}=\theta_{p}^{\circ n}\left(\mathfrak{z}\right)$, iterating (6.63) $\tau$ times gives:
(6.64) $X_{H}\left(\mathfrak{x}\right)=H_{\mathbf{j}}\left(X_{H}\left(\theta^{\circ\tau}\left(\mathfrak{x}\right)\right)\right)=H_{\mathbf{j}}\left(X_{H}\left(\mathfrak{x}\right)\right)$
which shows that $X_{H}\left(\mathfrak{x}\right)$ is fixed by $H_{\mathbf{j}}$. By (ii), using the notion of correctness and the arguments made with it in [14], it follows that for a string $\mathbf{j}$ of length $n$, the extension of $H_{\mathbf{j}}$ to $\mathcal{O}_{K_{v}}$ (which exists, since $v\left(\Lambda\right)\geq 1$), $H_{\mathbf{j}}\left(w\right)=w$ occurs if and only if $H_{\mathbf{j}}\left(w\right)=H^{\circ n}\left(w\right)$ and $H^{\circ n}\left(w\right)=w$. This makes $w=X_{H}\left(\mathfrak{x}\right)$ into a periodic point of $H$’s extension to $\mathcal{O}_{K_{v}}$. Moreover, we have that $X_{H}\left(\mathfrak{z}\right)=z$ lies in the forward orbit of $X_{H}\left(\mathfrak{x}\right)=w$ under $H$. By the integrality property of $H$, this forces $w$ to be in $\mathcal{O}_{K}$ since $z$ is, which shows that $z$ and $w$ are both periodic points of $H$ in $\mathcal{O}_{K}$. $\checkmark$
[…]
Q.E.D.
7. $p$-Adic Fourier Analysis and the $\ell$-adic Characteristic Function of $X_{H}$
Here, we introduce the Haar measure on $\mathbb{Z}_{p}$ and Pontryagin duality, the two cornerstones of $p$-adic Fourier analysis. The Haar measure is the natural generalization of the Lebesgue measure to locally compact abelian groups such as $\mathbb{Z}_{p}$, and, like with the Lebesgue integral on $\mathbb{R}$, allows us to do Fourier analysis with functions on $\mathbb{Z}_{p}$.
Definition 18.
Letting $p\geq 2$, the sets:
(7.1) $k+p^{n}\mathbb{Z}_{p}\overset{\textrm{def}}{=}\left\{k+p^{n}\mathfrak{z}:\mathfrak{z}\in\mathbb{Z}_{p}\right\},\textrm{ }\forall n\geq 0,\textrm{ }\forall k\in\left\{0,\ldots,p^{n}-1\right\}$
are both compact and open (“compact-open”) in $\mathbb{Z}_{p}$, and form a topological base for $\mathbb{Z}_{p}$. As such, they generate the Borel $\sigma$-algebra on $\mathbb{Z}_{p}$. The $p$-adic Haar measure $d\mathfrak{z}$ is defined on the Borel $\sigma$-algebra by the rule:
(7.2) $\int_{k+p^{n}\mathbb{Z}_{p}}d\mathfrak{z}\overset{\textrm{def}}{=}\frac{1}{p^{n}},\textrm{ }\forall n\geq 0,\textrm{ }\forall k\in\left\{0,\ldots,p^{n}-1\right\}$
This is the unique translation-invariant probability measure on $\mathbb{Z}_{p}$. For $r\in\left[1,\infty\right]$, $L^{r}\left(\mathbb{Z}_{p},\mathbb{C}\right)$ is then defined as the Banach space of functions $f:\mathbb{Z}_{p}\rightarrow\mathbb{C}$ under the usual norm:
(7.3) $\left\|f\right\|_{L^{r}\left(\mathbb{Z}_{p},\mathbb{C}\right)}\overset{\textrm{def}}{=}\begin{cases}\left(\int_{\mathbb{Z}_{p}}\left|f\left(\mathfrak{z}\right)\right|^{r}d\mathfrak{z}\right)^{1/r}&\textrm{if }r<\infty\\ \textrm{esssup}_{\mathfrak{z}\in\mathbb{Z}_{p}}\left|f\left(\mathfrak{z}\right)\right|&\textrm{if }r=\infty\end{cases}$
with the standard caveat that each function is actually an equivalence class of functions whose differences vanish almost everywhere with respect to $d\mathfrak{z}$.
Next, given $n\geq 0$ and $k\in\left\{0,\ldots,p^{n}-1\right\}$, we write:
(7.4) $\left[\mathfrak{z}\overset{p^{n}}{\equiv}k\right]\overset{\textrm{def}}{=}\begin{cases}1&\textrm{if }\mathfrak{z}\in k+p^{n}\mathbb{Z}_{p}\\ 0&\textrm{else}\end{cases},\textrm{ }\forall\mathfrak{z}\in\mathbb{Z}_{p}$
This function is said to be locally constant mod $p^{n}$, because its value depends only on the value of $\mathfrak{z}$ mod $p^{n}$. More generally, a function of a $p$-adic integer variable is said to be locally constant if it is locally constant mod $p^{n}$ for some $n\geq 0$.
A $p$-adic Schwartz-Bruhat (SB) function is a linear combination of finitely many indicator functions of this type. Every SB function is locally constant and Haar integrable, and every locally constant function is SB and Haar integrable. We write $\mathcal{S}\left(\mathbb{Z}_{p},\mathbb{C}\right)$ to denote the space of SB functions. This is an algebra under point-wise addition and multiplication, and is dense in $L^{r}\left(\mathbb{Z}_{p},\mathbb{C}\right)$ for all $r$. SB functions are to the classical Schwartz functions of Fourier analysis on euclidean space to what the Haar measure is to the Lebesgue.
Finally, although we will not need it for this paper, the Haar measure naturally extends from $\mathbb{Z}_{p}$ to the field $\mathbb{Q}_{p}$. Clopen balls of the form $\mathfrak{x}+p^{n}\mathbb{Z}_{p}$ for $\mathfrak{x}\in\mathbb{Q}_{p}$ and $n\in\mathbb{Z}$ form a basis for the topology of $\mathbb{Q}_{p}$. The $p$-adic Haar measure of $\mathfrak{x}+p^{n}\mathbb{Z}_{p}$ is:
(7.5) $\int_{\mathfrak{x}+p^{n}\mathbb{Z}_{p}}d\mathfrak{z}\overset{\textrm{def}}{=}p^{-n},\textrm{ }\forall\mathfrak{x}\in\mathbb{Q}_{p},\textrm{ }\forall n\in\mathbb{Z}$
Note that this makes intuitive sense, as $p^{n}\mathbb{Z}_{p}$ becomes a $p$-adically smaller neighborhood of 0 as $n\rightarrow+\infty$, while it becomes a $p$-adically larger neighborhood of 0 as $n\rightarrow$-$\infty$.
Remark 16.
Henceforth, unless stated otherwise, “almost everywhere” , “measure”, “measurable”, “integrable”, etc., are all meant with respect to the Haar measure.
Remark 17.
If $f:\mathbb{Z}_{p}\rightarrow\mathbb{C}$ is continuous (such as if $f$ is locally constant), it can be shown that:
(7.6) $\int_{\mathbb{Z}_{p}}f\left(\mathfrak{z}\right)d\mathfrak{z}=\lim_{N\rightarrow\infty}\frac{1}{p^{N}}\sum_{n=0}^{p^{N}-1}f\left(n\right)$
Moreover, given any $f:\mathbb{Z}_{p}\rightarrow\mathbb{C}$, the convergence of the limit (7.6) then implies that $f$ is Haar integrable, and that its integral is given by (7.6). See [15] for details.
Remark 18.
Because $\mathbb{Z}_{p}$ is a finite measure space, for all $r\geq 1$ and all measurable $f:\mathbb{Z}_{p}\rightarrow\mathbb{C}$, Hölder’s inequality gives:
(7.7) $\left\|f\right\|_{L^{r}\left(\mathbb{Z}_{p},\mathbb{C}\right)}\leq\left\|f\right\|_{L^{qr}\left(\mathbb{Z}_{p},\mathbb{C}\right)},\textrm{ }\forall r,q\geq 1$
Thus, $f\in L^{r_{1}}$ implies $f\in L^{r_{2}}$ for all $r_{1},r_{2}\in\left[1,\infty\right]$ with $r_{2}\leq r_{1}$. So, $f\in L^{\infty}\left(\mathbb{Z}_{p},\mathbb{C}\right)$ implies $f\in L^{r}\left(\mathbb{Z}_{p},\mathbb{C}\right)$ for all $r\in\left[1,\infty\right]$.
We need the following elementary integration formulae:
Proposition 11 (Decomposition formula).
Let $n\geq 0$. Then, for all measurable $f:\mathbb{Z}_{p}\rightarrow\mathbb{C}$:
(7.8) $\int_{\mathbb{Z}_{p}}f\left(\mathfrak{z}\right)d\mathfrak{z}=\sum_{k=0}^{p^{n}-1}\int_{k+p^{n}\mathbb{Z}_{p}}f\left(\mathfrak{z}\right)d\mathfrak{z}=\sum_{k=0}^{p^{n}-1}\int_{\mathbb{Z}_{p}}\left[\mathfrak{z}\overset{p^{n}}{\equiv}k\right]f\left(\mathfrak{z}\right)d\mathfrak{z}$
as well as the change of variables formula:
[…]
Using the change of variables formula, we also have:
(7.10) $\int_{\mathbb{Z}_{p}}f\left(\mathfrak{z}\right)d\mathfrak{z}=\frac{1}{p^{n}}\sum_{k=0}^{p^{n}-1}\int_{\mathbb{Z}_{p}}f\left(p^{n}\mathfrak{z}+k\right)d\mathfrak{z}$
Next, we have to introduce Pontryagin duality. This is the formalism which allows one to perform Fourier analysis for functions out of a locally compact abelian group (LCAG). LCAGs are abelian groups that are equipped with a topological structure with respect to which the operations of addition and negation are continuous. $\mathbb{R}$, $\mathbb{C}$, $\mathbb{Z}_{p}$, $\mathbb{Z}$, $\mathbb{Q}_{p}$, and $\hat{\mathbb{Z}}_{p}$ are all examples of LCAGs, though for our purposes, we will only need to work with $\mathbb{Z}_{p}$ and $\hat{\mathbb{Z}}_{p}$. We borrow the “frequency” terminology from Tao’s blog notes on Fourier Analysis [23].
Definition 19.
[…]
(7.11) $\chi\left(x+y\right)$ $=\chi\left(x\right)\chi\left(y\right),\textrm{ }\forall x,y\in G$
(7.12) $\left|\chi\left(z\right)\right|$ $=1,\textrm{ }\forall z\in G$
The trivial character is the unique character satisfying $\chi\left(x\right)=1$ for all $x\in G$. A frequency is a continuous group homomorphism $\xi:\left(G,+\right)\rightarrow\left(\mathbb{R}/\mathbb{Z},+\right)$ so that:
(7.13) $\xi\left(x+y\right)=\xi\left(x\right)+\xi\left(y\right),\textrm{ }\forall x,y\in G$
The trivial frequency is the unique frequency satisfying $\xi\left(x\right)=0$ for all $x\in G$.
Every unitary character $\chi$ can be uniquely written as:
(7.14) $\chi\left(x\right)=e^{-2\pi i\xi\left(x\right)},\textrm{ }\forall x\in G$
for some frequency $\xi$. The Pontryagin dual of $G$, denoted $\hat{G}$, is the set of all unitary characters, made into a group by pointwise multiplication; we can also define it as the set of all frequencies, made into a group by pointwise addition. These are the multiplicative and additive realizations of the same object. Pontryagin duality is the observation that the pairing:
(7.15) $\left(\xi,x\right)\in\hat{G}\times G\mapsto\xi\left(x\right)\in\mathbb{R}/\mathbb{Z}$
is a $\mathbb{Z}$-bilinear map called the duality bracket, and that this shows that $\hat{G}$ is also an LCAG. The Fourier transform then sends functions $X:G\rightarrow\mathbb{C}$ to functions $\hat{X}:\hat{G}\rightarrow\mathbb{C}$ defined by:
(7.16) $\hat{X}\left(\xi\right)\overset{\textrm{def}}{=}\int_{G}X\left(x\right)e^{-2\pi i\xi\left(x\right)}dx$
where $dx$ is the Haar measure of $G$, provided the integral exists. The Inverse Fourier Transform sends functions $\hat{X}:\hat{G}\rightarrow\mathbb{C}$ to functions $X:G\rightarrow\mathbb{C}$:
(7.17) $X\left(x\right)=\int_{\hat{G}}\hat{X}\left(\xi\right)e^{2\pi i\xi\left(x\right)}d\xi$
where $d\xi$ is the dual measure, the Haar measure on $\hat{G}$, provided the integral exists.
In our case, it can be shown that, as an additive group, the Pontryagin dual $\hat{\mathbb{Z}}_{p}$ is isomorphic to $\mathbb{Z}\left[1/p\right]/\mathbb{Z}$, the group of all $p$-power denominator rational numbers in $\left[0,1\right)$ with addition mod $1$ as the group operation. The dual measure on $\hat{\mathbb{Z}}_{p}$ is just the counting measure. The duality bracket is given explicitly by the $p$-adic fractional part, $\left\{\cdot\right\}_{p}:\mathbb{Q}_{p}\rightarrow\hat{\mathbb{Z}}_{p}$. If $p$ is a prime number, this is defined by:
(7.18) $\left\{\mathfrak{x}\right\}_{p}=\begin{cases}0&\textrm{if }n_{0}\geq 0\\ \sum_{n=n_{0}}^{-1}d_{n}p^{n}&\textrm{if }n_{0}\leq-1\end{cases}$
where:
(7.19) $\mathfrak{x}=\sum_{n=n_{0}}^{\infty}d_{n}p^{n}$
is the Hensel series representation of $\mathfrak{x}\in\mathbb{Q}_{p}$. If $p$ is composite, $\left\{\cdot\right\}_{p}$ is defined by:
(7.20) $\left\{\mathfrak{x}\right\}_{p}\overset{\textrm{def}}{=}\sum_{\ell\mid p}\left\{\mathfrak{x}\right\}_{\ell}$
where the sum is taken over all prime divisors $\ell$ of $p$.
Thus, for example, given $t\in\hat{\mathbb{Z}}_{p}$ and $\mathfrak{z}\in\mathbb{Z}_{p}$:
(7.21) $e^{2\pi i\left\{t\mathfrak{z}\right\}_{p}}=\begin{cases}1&\textrm{if }t=0\\ e^{2\pi it\left[\mathfrak{z}\right]_{\left|t\right|_{p}}}&\textrm{else}\end{cases}$
where $\left[\mathfrak{z}\right]_{\left|t\right|_{p}}$ is the projection of $\mathfrak{z}$ mod the power of $p$ in $t$’s denominator. Thus, if $p$ is odd:
(7.22) $e^{2\pi i\left\{2\mathfrak{z}/p^{3}\right\}_{p}}=e^{2\pi i\left(2\left[\mathfrak{z}\right]_{p^{3}}/p^{3}\right)}$
Moreover, we have that:
(7.23) $e^{2\pi i\left\{tn\right\}_{p}}=e^{2\pi itn},\textrm{ }\forall n\in\mathbb{Z},\textrm{ }\forall t\in\hat{\mathbb{Z}}_{p}$
and that:
(7.24) $e^{2\pi i\left\{\frac{t}{n}\right\}_{p}}=e^{2\pi it\left[n\right]_{\left|t\right|_{p}}^{-1}},\forall t\in\hat{\mathbb{Z}}_{p}\backslash\left\{0\right\},\textrm{ }\forall n\in\mathbb{Z}:\gcd\left(n,p\right)=1$
where $\left[n\right]_{\left|t\right|_{p}}^{-1}$ is the unique integer in $\left\{0,\ldots,\left|t\right|_{p}-1\right\}$ whose product with $n$ is congruent to $1$ mod $\left|t\right|_{p}$.
Given any function $X\in L^{1}\left(\mathbb{Z}_{p},\mathbb{C}\right)$, the Fourier transform of $X$, denoted $\hat{X}:\hat{\mathbb{Z}}_{p}\rightarrow\mathbb{C}$ is the function defined by the integral:
(7.25) $\hat{X}\left(t\right)\overset{\textrm{def}}{=}\int_{\mathbb{Z}_{p}}X\left(\mathfrak{z}\right)e^{-2\pi i\left\{t\mathfrak{z}\right\}_{p}}d\mathfrak{z}$
Given a function $\hat{Y}:\hat{\mathbb{Z}}_{p}\rightarrow\mathbb{C}$, the Fourier series generated by $\hat{Y}$ is:
(7.26) $\mathfrak{z}\mapsto\sum_{t\in\hat{\mathbb{Z}}_{p}}\hat{Y}\left(t\right)e^{2\pi i\left\{t\mathfrak{z}\right\}_{p}}$
where we write $\sum_{t\in\hat{\mathbb{Z}}_{p}}\hat{X}\left(t\right)$ to denote the limit:
(7.27) $\sum_{t\in\hat{\mathbb{Z}}_{p}}\hat{X}\left(t\right)\overset{\textrm{def}}{=}\lim_{N\rightarrow\infty}\sum_{\left|t\right|_{p}\leq p^{N}}\hat{X}\left(t\right)\overset{\textrm{def}}{=}\lim_{N\rightarrow\infty}\sum_{k=0}^{p^{N}-1}\hat{X}\left(\frac{k}{p^{N}}\right)$
Remark 19.
[…]
(7.28) $\int_{G}f\left(x\right)g\left(x\right)dx=\int_{\hat{G}}\hat{f}\left(\xi\right)\hat{g}\left(-\xi\right)d\xi$
valid for all $f$ and $g$ for which either the left-hand or right-hand sides converge. As with the classical Parseval-Plancherel Identity, we can use (7.28) to define measures and distributions by specifying how they act in terms of frequencies; ex:
(7.29) $\int_{G}f\left(x\right)d\mu\left(x\right)=\int_{\hat{G}}\hat{f}\left(\xi\right)\hat{\mu}\left(-\xi\right)d\xi$
where
(7.30) $\hat{\mu}\left(\xi\right)=\int_{G}e^{-2\pi i\xi\left(x\right)}d\mu\left(x\right)$
is the Fourier-Stieltjes transform of the measure $d\mu$ on $G$.
Remark 20.
The theory of integration and Fourier analysis we sketched above is extremely robust, in that it can be transferred almost verbatim to the case of functions $\mathbb{Z}_{p}\rightarrow R$, where $R$ is any commutative, unital, metrically complete normed ring so that both $R$ and $R$’s residue field (in the case where $R$ is a non-archimedean local ring) have characteristic either 0 or coprime to $p$. See [15] for a comprehensive exposition of this material.
By Lemma 1, the condition $\rho_{H,\ell}<1$ guarantees that $X_{H}:\mathbb{Z}_{p}\rightarrow K_{\ell}$ is measurable. As such, we can view $X_{H}$ as a $K_{\ell}$-valued random variable whose associated sample space is the Borel $\sigma$-algebra of $\mathbb{Z}_{p}$. Given any measurable subset $A$ of $K_{\ell}$, the probability $\mathbb{F}\left(X_{H}\in A\right)$ would then be the Haar probability measure of the subset of $\mathbb{Z}_{p}$ on which $X_{H}\in A$.
In viewing $X_{H}$ as a random variables, it is natural to want to consider their characteristic function.
Definition 20.
Let $K$ be a number field of dimension $d$ over $\mathbb{Q}$, and let $\ell$ be a place of $K$. By algebraic number theory, there exists a unique place $q$ of $\mathbb{Q}$ so that $\ell$ lies over $q$. Letting $\mathbb{Q}_{q}$ denote the completion of $\mathbb{Q}$ with respect to $q$, given a $p$-Hydra $H:\mathcal{O}_{K}\rightarrow\mathcal{O}_{K}$, the $\ell$-adic characteristic function of $X_{H}$, denoted $\hat{\mu}_{H,\ell}$, is the function $\hat{\mu}_{H,\ell}:\hat{\mathcal{O}}_{K_{\ell}}\rightarrow\mathbb{C}$ defined by:
(7.31) $\hat{\mu}_{H,\ell}\left(t\right)\overset{\textrm{def}}{=}\int_{\mathbb{Z}_{p}}e^{-2\pi i\left\{\textrm{Tr}_{K_{\ell}/\mathbb{Q}_{q}}\left(tX_{H}\left(\mathfrak{z}\right)\right)\right\}_{q}}d\mathfrak{z},\textrm{ }\forall t\in\hat{\mathcal{O}}_{K_{\ell}}$
Here, when $\ell$ is non-archimedean, $\hat{\mathcal{O}}_{K_{\ell}}$ is defined to be $K_{\ell}/\mathcal{O}_{K_{\ell}}$, the pontryagin dual of $\mathcal{O}_{K_{\ell}}$, the ring of integers of the $\ell$-adic completion of $K$. When $\ell$ is archimedean, $\hat{\mathcal{O}}_{K_{\ell}}$ is defined to be $\mathbb{R}$. $\textrm{Tr}_{K_{\ell}/\mathbb{Q}_{q}}$ is the field trace of the extension $K_{\ell}/\mathbb{Q}_{q}$, and $\left\{\cdot\right\}_{q}$ is defined to be the $q$-adic fractional part $\left\{\cdot\right\}_{q}:\mathbb{Q}_{q}\rightarrow\hat{\mathbb{Z}}_{q}\subseteq\mathbb{R}/\mathbb{Z}$ when $\ell$ is non-archimedean and is defined to be the identity map $\left\{\cdot\right\}_{q}:\mathbb{R}\rightarrow\mathbb{R}$ when $\ell$ is archimedean. $X_{H}$, of course, is the numen of $H$. We write $e_{\ell}:K_{\ell}\rightarrow\mathbb{T}$ to denote:
(7.32) $e_{\ell}\left(\mathfrak{x}\right)=\exp\left(2\pi i\left\{\textrm{Tr}_{K_{\ell}/\mathbb{Q}_{q}}\left(\mathfrak{x}\right)\right\}\right)_{q},\textrm{ }\forall\mathfrak{x}\in K_{\ell}$
With this notation, we can write $\hat{\mu}_{H,\ell}$ as:
(7.33) $\hat{\mu}_{H,\ell}\left(t\right)=\int_{\mathbb{Z}_{p}}e_{\ell}\left(-tX_{H}\left(\mathfrak{z}\right)\right)d\mathfrak{z}$
Remark 21.
Note that the measurability of $X_{H}$ guaranteed by $\rho_{H,\ell}$ ensures that the integral defining $\hat{\mu}_{H,\ell}$ is well-defined and absolutely convergent with respect to $t$.
Proposition 12.
We have:
(7.34) $\hat{\mu}_{H,\ell}\left(t\right)=\frac{1}{p}\sum_{j=0}^{p-1}e_{\ell}\left(-c_{j}t\right)\hat{\mu}_{H,\ell}\left(r_{j}t\right)$
Using these functional equations, one can show that $\hat{\mu}_{H,\ell}$ is the Fourier transform of the unique self-similar probability measure $\mu_{H,\ell}$ associated to the IFS generated by the $H_{j}$s. For this, note that by Pontryagin duality, given a function $f:K_{\ell}\rightarrow\mathbb{C}$, its Fourier transform $\hat{f}:K_{\ell}\rightarrow\mathbb{C}$ is:
$\hat{f}\left(w\right)=\int_{K_{\ell}}f\left(z\right)e_{\ell}\left(-wz\right)dz,\textrm{ }\forall w\in K_{\ell}$
where $dz$ is the Haar measure on $K_{\ell}$, with $dz$ being normalized to be the Haar probability measure on $\mathcal{O}_{K_{\ell}}$ if $\ell$ is non-archimedean. We also need:
Definition 21.
Let $W\left(K_{\ell},\mathbb{C}\right)$ denote the set of all functions $\phi:K_{\ell}\rightarrow\mathbb{C}$ so that:
(7.35) $\phi\left(z\right)=\int_{K_{\ell}}\hat{\phi}\left(w\right)e_{\ell}\left(wz\right)dw,\textrm{ }\forall z\in K_{\ell}$
for some $\hat{\phi}\in L^{1}\left(K_{\ell},\mathbb{C}\right)$. This is the Wiener algebra of functions $\phi:K_{\ell}\rightarrow\mathbb{C}$ with absolutely integrable Fourier transforms, and is a Banach algebra with respect to pointwise multiplication under the norm:
(7.36) $\left\|\phi\right\|_{W\left(K_{\ell},\mathbb{C}\right)}\overset{\textrm{def}}{=}\left\|\hat{\phi}\right\|_{L^{1}\left(K_{\ell},\mathbb{C}\right)}\overset{\textrm{def}}{=}\int_{K_{\ell}}\left|\hat{\phi}\left(w\right)\right|dw$
When $\ell$ is a non-archimedean place, we write $W\left(\mathcal{O}_{K_{\ell}},\mathbb{C}\right)$ to denote the subspace of $W\left(K_{\ell},\mathbb{C}\right)$ consisting of all functions in $W\left(K_{\ell},\mathbb{C}\right)$ whose supports lie in $\mathcal{O}_{K_{\ell}}$.
The connection between self-similar measures ([20] is a good survey article on the subject) and our formalism is then as follows:
Proposition 13.
Let $d\mu_{H,\ell}$ denote the continuous linear functional $W\left(K_{\ell},\mathbb{C}\right)\rightarrow\mathbb{C}$ defined by:
(7.37) $\int_{K_{\ell}}f\left(z\right)d\mu_{H,\ell}\left(z\right)\overset{\textrm{def}}{=}\int_{K_{\ell}}\hat{f}\left(t\right)\hat{\mu}_{H,\ell}\left(-t\right)dt,\textrm{ }\forall f\in W\left(K_{\ell},\mathbb{C}\right)$
That is:
(7.38) $\hat{\mu}_{H,\ell}\left(t\right)=\int_{K_{\ell}}e_{\ell}\left(-tz\right)d\mu_{H}\left(z\right)$
Then, $d\mu_{H,\ell}$ extends to a probability measure on $K_{\ell}$, and is the unique self-similar Borel probability measure on $K_{\ell}$ satisfying:
(7.39) $d\mu_{H,\ell}=\frac{1}{p}\sum_{j=0}^{p-1}H_{j,*}\left\{d\mu_{H,\ell}\right\}=\frac{1}{p}\sum_{j=0}^{p-1}d\mu_{H,\ell}\circ H_{j}^{-1}$
i.e.:
(7.40) $\int_{K_{\ell}}f\left(z\right)d\mu_{H,\ell}\left(z\right)=\frac{1}{p}\sum_{j=0}^{p-1}\int_{K_{\ell}}f\left(H_{j}\left(z\right)\right)d\mu_{H,\ell}\left(z\right)$
In particular, we have that:
(7.41) $\int_{K_{\ell}}f\left(z\right)d\mu_{H,\ell}\left(z\right)=\int_{\mathbb{Z}_{p}}f\left(X_{H}\left(\mathfrak{z}\right)\right)d\mathfrak{z},\textrm{ }\forall f\in L^{1}\left(K_{\ell},\mathbb{C}\right)$
Proof: Self-similarity is straightforward:
$\int_{K_{\ell}}f\left(z\right)d\mu_{H}\left(z\right)$ $=\int_{K_{\ell}}\hat{f}\left(t\right)\hat{\mu}_{H,\ell}\left(-t\right)dt$
[…]
$\left(\textrm{Def. of }d\mu_{H,\ell}\right);$ $=\frac{1}{p}\sum_{j=0}^{p-1}\int_{K_{\ell}}\hat{f}\left(t\right)e_{\ell}\left(c_{j}t\right)\left(\int_{V_{\ell}}e_{\ell}\left(tr_{j}z\right)d\mu_{H,\ell}\left(z\right)\right)dt$
$\left(\textrm{Fubini}\right);$ $=\frac{1}{p}\sum_{j=0}^{p-1}\int_{K_{\ell}}\left(\int_{K_{\ell}}\hat{f}\left(t\right)e_{\ell}\left(t\left(r_{j}z+c_{j}\right)\right)dt\right)d\mu_{H,\ell}\left(z\right)$
$\left(f\in W\left(K_{\ell},\mathbb{C}\right)\right);$ $=\frac{1}{p}\sum_{j=0}^{p-1}\int_{K_{\ell}}f\left(r_{j}z+c_{j}\right)d\mu_{H,\ell}\left(z\right)$
$=\frac{1}{p}\sum_{j=0}^{p-1}\int_{K_{\ell}}f\left(H_{j}\left(z\right)\right)d\mu_{H,\ell}\left(z\right)$
As for extendibility, the construction of $d\mu_{H,\ell}$ guarantees that:
(7.42) $\int_{K_{\ell}}f\left(z\right)d\mu_{H,\ell}\left(z\right)=\int_{\mathbb{Z}_{p}}f\left(X_{H}\left(\mathfrak{z}\right)\right)d\mathfrak{z},\textrm{ }\forall f\in W\left(K_{\ell},\mathbb{C}\right)$
This follows from the fact that:
$\int_{K_{\ell}}f\left(z\right)d\mu_{H,\ell}\left(z\right)$ $=\int_{K_{\ell}}\hat{f}\left(w\right)\hat{\mu}_{H,\ell}\left(w\right)dw$
$\left(\textrm{def. of }\hat{\mu}_{H,\ell}\right);$ $=\int_{K_{\ell}}\hat{f}\left(w\right)\left(\int_{\mathbb{Z}_{p}}e_{\ell}\left(-wX_{H}\left(\mathfrak{z}\right)\right)d\mathfrak{z}\right)dw$
$\left(\textrm{swap integrals}\right);$ $\overset{!}{=}\int_{\mathbb{Z}_{p}}\left(\int_{K_{\ell}}\hat{f}\left(w\right)e_{\ell}\left(-wX_{H}\left(\mathfrak{z}\right)\right)dw\right)d\mathfrak{z}$
$\left(f\in W\left(K_{\ell},\mathbb{C}\right)\right);$ $=\int_{\mathbb{Z}_{p}}f\left(X_{H}\left(\mathfrak{z}\right)\right)d\mathfrak{z}$
where the interchange of integrals at the step marked (!) is justified by the $L^{1}$ convergence of the $dw$ integral. Because of this, we see that:
$\int_{K_{\ell}}f\left(z\right)d\mu_{H,\ell}\left(z\right)\geq 0,\textrm{ }\forall f\in W\left(K_{\ell},\mathbb{C}\right):f\left(z\right)\geq 0\textrm{ }\forall z\in\mathbb{C}$
and hence, that $d\mu_{H,\ell}$ is a positive linear functional. From this and the containment of $W\left(K_{\ell},\mathbb{C}\right)$ in the space $C_{0}\left(K_{\ell},\mathbb{C}\right)$ of continuous functions $K_{\ell}\rightarrow\mathbb{C}$ that decay to 0 as their input variable tends to $\infty$ in magnitude in $K_{\ell}$, it follows by the Riesz-Markov-Katukani Representation Theorem that $d\mu_{H,\ell}$ extends to a Borel probability measure on $K_{\ell}$. The self-similarity condition proved above guarantees uniqueness. The aforementioned extension property then lets us extend (7.42) to all $f\in L^{1}$, thus demonstrating (7.41).
Q.E.D.
In the case where $\ell$ is non-archimedean, we can do more. Since $q$ is the unique prime of $\mathbb{Q}$ lying under $\ell$, it follows that $K_{\ell}$ is isomorphic to a finite-degree extension $F$ of $\mathbb{Q}_{q}$. Moreover, $\left|\cdot\right|_{q}$ has a unique extension to an absolute value on $F$, and this extension is equal to the absolute value $\left|\cdot\right|_{\ell}$ on $K_{\ell}$. Thus, we may speak meaningfully about the $q$-adic absolute values of elements of $K_{\ell}$. To that end, we have the following identities, which are useful for computations.
Proposition 14.
Let $q$ lie under $\ell$, and let $\pi_{\ell}$ be a uniformizer of $\mathcal{O}_{K_{\ell}}$, and let $e_{K,\ell}$ be the ramification index of $K_{\ell}$ over $\mathbb{Q}_{q}$. Then, in $\hat{\mathcal{O}}_{K_{\ell}}=K_{\ell}/\mathcal{O}_{K_{\ell}}$, for any $n\in\mathbb{N}_{0}$, we have:
(7.43) $\left|\left\{t\in\hat{\mathcal{O}}_{K_{\ell}}:\left|t\right|_{\ell}\leq q^{n/e_{K,\ell}}\right\}\right|=\left(q^{n/e_{K,\ell}}\right)^{\left[K_{\ell}:\mathbb{Q}_{q}\right]}$
Proposition 15.
Let $m,n$ be non-negative integers with $n\geq m$, let $A$ be an abelian group, and consider functions $\hat{f},\hat{g}:\hat{\mathcal{O}}_{K_{\ell}}\longrightarrow A$. Then:
(7.44) $\sum_{\left|t\right|_{\ell}\leq q^{n/e_{K,\ell}}}\hat{f}\left(\pi_{\ell}^{m}\right)\hat{g}\left(t\right)=\sum_{\left|t\right|_{\ell}\leq q^{\left(n-m\right)/e_{K,\ell}}}\left(\sum_{\left|t\right|_{\ell}\leq q^{m/e_{K,\ell}}}\hat{f}\left(s+\pi_{\ell}^{-m}t\right)\right)\hat{g}\left(t\right)$
Finally, using Fourier inversion, we can express the $\ell$-adic probability distribution of $X_{H}$ in terms of $\hat{\mu}_{H,\ell}$.
Proposition 16.
Suppose that $\ell$ is a non-archimedean place for which $\rho_{H,\ell}<1$ and $\max_{0\leq j<p}\left\|r_{j}\right\|_{\ell}\leq 1$. Then:
(7.45) $\mathbb{P}\left(X_{H}\overset{\pi_{\ell}^{n}}{\equiv}\mathfrak{w}\right)=q^{-n/e_{K,\ell}}\sum_{\left|t\right|_{\ell}\leq q^{n/e_{K,\ell}}}\hat{\mu}_{H,\ell}\left(t\right)e_{\ell}\left(t\mathfrak{w}\right),\textrm{ }\forall\mathfrak{w}\in\pi_{\ell}^{-B_{H,\ell}}\mathcal{O}_{K_{\ell}}$
for all $n\in\mathbb{N}_{0}$, where $B_{H,\ell}$ is the integer:
(7.46) $B_{H,\ell}\overset{\textrm{def}}{=}e_{K,\ell}\max_{0\leq j<p}\log_{q}\left|c_{j}\right|_{\ell}$
where $\log_{q}x=\left(\ln x\right)/\ln q$.
Remark 22.
The probability on the left can also be written as:
(7.47) $\mathbb{P}\left(X_{H}\overset{q^{n/e_{K,\ell}}}{\equiv}\mathfrak{w}\right)$
Proof: By (6.25) from Lemma 1, the hypotheses on the $\ell$-adic norms of the $r_{j}$s implies that $X_{H}$ is $\ell$-adically bounded by:
(7.48) $\max_{0\leq j<p}\left|c_{j}\right|_{\ell}$
for all $\mathfrak{z}$ for which $X_{H}\left(\mathfrak{z}\right)$ is $\ell$-adically convergent. Since the $\ell$-adic absolute value takes non-zero values of the form $q^{n/e_{K,\ell}}$ for $n\in\mathbb{Z}$, it follows that there is some constant $B_{H,\ell}\in\mathbb{Z}$ so that:
$\max_{0\leq j<p}\left|c_{j}\right|_{\ell}=q^{B_{H,\ell}/e_{K,\ell}}=\left|\pi_{\ell}^{B_{H,\ell}}\right|_{q}$
By character orthogonality, using (14), we have:
$\sum_{\left|t\right|_{\ell}\leq q^{n/e_{K,\ell}}}e_{\ell}\left(t\left(\mathfrak{y}-\pi_{\ell}^{B_{H,\ell}}\mathfrak{x}\right)\right)=\left(q^{n/e_{K,\ell}}\right)^{\left[K_{\ell}:\mathbb{Q}_{q}\right]}\left[\pi_{\ell}^{B_{H,\ell}}\mathfrak{x}\overset{q^{n/e_{K,\ell}}}{\equiv}\mathfrak{y}\right]$
for all $\mathfrak{x}\in K_{\ell}$ with $\left|\mathfrak{x}\right|_{\ell}\leq\pi_{\ell}^{B_{H,\ell}}$ and all $\mathfrak{y}\in\mathcal{O}_{K_{\ell}}$. The congruence in the Iverson bracket is well-defined because the bound on $\mathfrak{x}$ implies $\pi_{\ell}^{B_{H,\ell}}\mathfrak{x}$ is in $\mathcal{O}_{K_{\ell}}$. Consequently, our bound on $X_{H}$ implies:
$\sum_{\left|t\right|_{\ell}\leq q^{n/e_{K,\ell}}}e_{\ell}\left(t\left(\mathfrak{y}-\pi_{\ell}^{B_{H,\ell}}X_{H}\left(\mathfrak{z}\right)\right)\right)=\left(q^{n/e_{K,\ell}}\right)^{\left[K_{\ell}:\mathbb{Q}_{q}\right]}\left[\pi_{\ell}^{B_{H,\ell}}X_{H}\left(\mathfrak{z}\right)\overset{q^{n/e_{K,\ell}}}{\equiv}\mathfrak{y}\right]$
for all $\mathfrak{y}\in\mathcal{O}_{K_{\ell}}$ and almost every $\mathfrak{z}\in\mathbb{Z}_{p}$. Integrating with respect to $\mathfrak{z}$ over $\mathbb{Z}_{p}$, we have that:
$\int_{\mathbb{Z}_{p}}\left[\pi_{\ell}^{B_{H,\ell}}X_{H}\left(\mathfrak{z}\right)\overset{q^{n/e_{K,\ell}}}{\equiv}\mathfrak{y}\right]d\mathfrak{z}=\mathbb{P}\left(\pi_{\ell}^{B_{H,\ell}}X_{H}\overset{q^{n/e_{K,\ell}}}{\equiv}\mathfrak{y}\right)$
Thus:
$\left(q^{n/e_{K,\ell}}\right)^{\left[K_{\ell}:\mathbb{Q}_{q}\right]}\mathbb{P}\left(\pi_{\ell}^{B_{H,\ell}}X_{H}\overset{q^{n/e_{K,\ell}}}{\equiv}\mathfrak{y}\right)$ $=\sum_{\left|t\right|_{\ell}\leq q^{n/e_{K,\ell}}}\int_{\mathbb{Z}_{p}}e_{\ell}\left(t\left(\mathfrak{y}-\pi_{\ell}^{B_{H,\ell}}X_{H}\left(\mathfrak{z}\right)\right)\right)d\mathfrak{z}$
$=\sum_{\left|t\right|_{\ell}\leq q^{n/e_{K,\ell}}}e_{\ell}\left(t\mathfrak{y}\right)\hat{\mu}_{H,\ell}\left(\pi_{\ell}^{B_{H,\ell}}t\right)$
Dividing by the constant on the left-hand side gives:
(7.49) $\mathbb{P}\left(\pi_{\ell}^{B_{H,\ell}}X_{H}\overset{q^{n/e_{K,\ell}}}{\equiv}\mathfrak{y}\right)=\left(q^{-n/e_{K,\ell}}\right)^{\left[K_{\ell}:\mathbb{Q}_{q}\right]}\sum_{\left|t\right|_{\ell}\leq q^{n/e_{K,\ell}}}\hat{\mu}_{H,\ell}\left(\pi_{\ell}^{B_{H,\ell}}t\right)e_{\ell}\left(t\mathfrak{y}\right)$
Here:
$\mathbb{P}\left(\pi_{\ell}^{B_{H,\ell}}X_{H}\overset{q^{n/e_{K,\ell}}}{\equiv}\mathfrak{y}\right)$ $=\mathbb{P}\left(\left|\pi_{\ell}^{B_{H,\ell}}X_{H}-\mathfrak{y}\right|_{q}\leq q^{-n/e_{K,\ell}}\right)$
$=\mathbb{P}\left(\left|X_{H}-\frac{\mathfrak{y}}{\pi_{\ell}^{B_{H,\ell}}}\right|_{q}\leq q^{\left(B_{H,\ell}-n\right)/e_{K,\ell}}\right)$
$=\mathbb{P}\left(X_{H}\overset{q^{\left(n-B_{H,\ell}\right)/e_{K,\ell}}}{\equiv}\frac{\mathfrak{y}}{q^{B_{H,\ell}/e_{K,\ell}}}\right)$
Since:
$\left\{\pi_{\ell}^{-B_{H,\ell}}\mathfrak{y}:\mathfrak{y}\in\mathcal{O}_{K_{\ell}}\right\}=\pi_{\ell}^{-B_{H,\ell}}\mathcal{O}_{K_{\ell}}$
This gives:
$\mathbb{P}\left(X_{H}\overset{q^{\left(n-B_{H,\ell}\right)/e_{K,\ell}}}{\equiv}\mathfrak{w}\right)=\left(q^{-n/e_{K,\ell}}\right)^{\left[K_{\ell}:\mathbb{Q}_{q}\right]}\sum_{\left|t\right|_{\ell}\leq q^{n/e_{K,\ell}}}\hat{\mu}_{H,\ell}\left(\pi_{\ell}^{B_{H,\ell}}t\right)e_{\ell}\left(\pi_{\ell}^{B_{H,\ell}}t\mathfrak{w}\right)$
for all $\mathfrak{w}\in\pi_{\ell}^{-B_{H,\ell}}\mathcal{O}_{K_{\ell}}$. Letting $n\geq B_{H,\ell}$, we can apply Proposition 15 to get:
$\mathbb{P}\left(X_{H}\overset{q^{\left(n-B_{H,\ell}\right)/e_{K,\ell}}}{\equiv}\mathfrak{w}\right)$ $=\left(q^{-n/e_{K,\ell}}\right)^{\left[K_{\ell}:\mathbb{Q}_{q}\right]}\sum_{\left|t\right|_{\ell}\leq q^{\left(n-B_{H,\ell}\right)/e_{K,\ell}}}\hat{\mu}_{H,\ell}\left(t\right)e_{\ell}\left(t\mathfrak{w}\right)\sum_{\left|s\right|_{q}\leq q^{B_{H,\ell}/e_{K,\ell}}}1$
$=q^{-\left(n-B_{H,\ell}\right)\left[K_{\ell}:\mathbb{Q}_{q}\right]/e_{K,\ell}}\sum_{\left|t\right|_{\ell}\leq q^{\left(n-B_{H,\ell}\right)/e_{K,\ell}}}\hat{\mu}_{H,\ell}\left(t\right)e_{\ell}\left(t\mathfrak{w}\right)$
Replacing $n$ with $n+B_{H,\ell}$ then gives the result.
Q.E.D.